### v1

# Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction

Abstract
The proliferation of diverse remote sensing data has spurred advancements in dense prediction tasks, yet significant challenges remain in handling data heterogeneity. Remote sensing imagery exhibits substantial variability across temporal, spectral, and spatial (TSS) dimensions, complicating unified data processing. Current deep learning models for dense prediction tasks, such as semantic segmentation and change detection, are typically tailored to specific input-output configurations. Consequently, variations in data dimensionality or task requirements often lead to significant performance degradation or model incompatibility, necessitating costly retraining or fine-tuning efforts for different application scenarios. This paper introduces the Temporal-Spectral-Spatial Unified Network (TSSUN), a novel architecture designed for unified representation and modeling of remote sensing data across diverse TSS characteristics and task types. TSSUN employs a Temporal-Spectral-Spatial Unified Strategy that leverages meta-information to decouple and standardize input representations from varied temporal, spectral, and spatial configurations, and similarly unifies output structures for different dense prediction tasks and class numbers. Furthermore, a Local-Global Window Attention mechanism is proposed to efficiently capture both local contextual details and global dependencies, enhancing the model’s adaptability and feature extraction capabilities. Extensive experiments on multiple datasets, encompassing tasks like building detection and land use/land cover classification with diverse TSS configurations, demonstrate that a single TSSUN model effectively adapts to heterogeneous inputs and unifies various dense prediction tasks. The proposed approach consistently achieves or surpasses state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications without requiring task-specific modifications.
Index Terms:
Remote Sensing, Dense Prediction, Unified Network, Change Detection, Semantic Segmentation, Deep learning
I INTRODUCTION
With the continuous advancement of remote sensing technologies and the increasing diversity of data acquisition methods [1], the field of remote sensing has entered a phase of rapid development [2]. Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7]. Remote sensing imagery exhibits high heterogeneity across three key dimensions—temporal, spectral, and spatial—posing significant challenges for unified processing due to variations in temporal length, number of spectral channels, and spatial resolution in practical applications [8].
From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, semantic change detection, and binary change detection. These tasks can be formally defined as follows: given a remote sensing image of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the spatial dimensions of the input and output, respectively. Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$. In semantic segmentation, the model performs multi-class land cover extraction from a single time-point image, corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points, corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points, corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$.
In recent years, deep learning methods, supported by the growing abundance of remote sensing data, have achieved notable success in dense prediction tasks, leading to the emergence of a wide range of high-performing models across all three task categories. However, existing models are typically designed for fixed input-output configurations, i.e., specific $(T_{1},C_{1},H_{1},W_{1})$ and $(T_{2},C_{2},H_{2},W_{2})$. Even slight variations in any of these dimensions can lead to significant performance degradation or complete incompatibility. In particular, changes in $T_{1}$, $C_{1}$, and $(H_{1},W_{1})$ reflect the diversity of remote sensing data along the Temporal- Spectral-Spatial (TSS) dimensions; variations in $T_{2}$ correspond to different task types, while differences in $C_{2}$ relate to varying classification requirements. Given that remote sensing applications often differ across these dimensions, models trained in one scenario typically struggle to generalize to others, requiring additional training or fine-tuning, which introduces considerable computational and time costs.
Figure 1: Illustration of Temporal-Spectral-Spatial Unified Network. TSSUN is capable of handling input and output with arbitrary temporal, spectral, and spatial dimensions, supports dense prediction for any number of classes, and unifies semantic segmentation, binary change detection and semantic change detection tasks.
To address these issues, this study propose a Temporal-Spectral-Spatial Unified Network (TSSUN) that enables unified representation and modeling of remote sensing data across the temporal, spectral, and spatial dimensions, as shown in Figure 1. TSSUN is highly flexible, capable of handling any combination of $(T_{1},C_{1},H_{1},W_{1})$ inputs and $(T_{2},C_{2},H_{2},W_{2})$ outputs. It supports inputs with varying temporal lengths, spectral bands, and spatial resolutions, and unifies all three dense prediction tasks. Moreover, it can produce outputs with an arbitrary number of semantic classes and maintains strong performance across diverse configurations.
Specifically, we introduce the Temporal-Spectral-Spatial Unified Strategy (TSSUS) to decouple and unify the input and output representations across the TSS dimensions. At the input stage, TSSUS leverages inherent spectral and spatial characteristics to achieve unified encoding of remote sensing imagery. Along the spectral dimension, it exploits the continuity across spectral bands and integrates spectral metadata to accomplish spectral harmonization. In parallel, along the spatial dimension, it harnesses the similarity among nearby regions, supplemented with spatial metadata, to produce a unified representation of spatial scales. By preserving the time dimension at the input, TSSUS is able to independently extract features from remote sensing images across distinct temporal instances. At the encoder–decoder junction, TSSUS facilitates feature-level fusion along the temporal axis, effectively mitigating interference from irrelevant information while adapting the output temporal length to suit the dense prediction task at hand [9, 10]. At the output stage, TSSUS further capitalizes on the structured nature of remote sensing dense prediction tasks in both spectral and spatial dimensions. In the spectral dimension, by modeling inter-class correlations, it enables prediction over an arbitrary number of classes. Meanwhile, in the spatial dimension, by exploiting the spatial structural consistency among neighboring regions, it significantly enhances the model’s generalization capability when outputting data at varying spatial resolutions.
To further improve the model’s ability to capture diverse TSS combinations, we design a Local-Global Window Attention (LGWA) mechanism. This module efficiently extracts local features using three overlapping window-based attention blocks with different shapes, followed by a global attention block that aggregates information at the global level. This design achieves a balance between computational efficiency and expressive power, enabling collaborative modeling of local and global features and significantly boosting the model’s performance in complex remote sensing tasks.
In summary, the main contributions of this work are as follows:
  1. 1.
    A Temporal-Spectral-Spatial Unified Network is proposed to enable unified modeling across temporal, spectral, and spatial dimensions for dense prediction in remote sensing. TSSUN supports arbitrary TSS input configurations and accommodates various tasks, including semantic segmentation, semantic change detection, and binary change detection with flexible output class settings.
  2. 2.
    A Local-Global Window Attention mechanism is designed to efficiently capture both local and global contextual features, improving performance across a wide range of remote sensing prediction tasks.
  3. 3.
    Extensive experiments are conducted on multiple datasets with diverse TSS configurations, including building detection and land use/land cover (LULC) classification. The results show that a single TSSUN model can adapt to heterogeneous inputs, unify multiple task types, and consistently achieve or exceed state-of-the-art (SOTA) performance.
II Related Works
II-A Task Unification in Remote Sensing Dense Prediction
Dense prediction tasks in remote sensing, particularly semantic segmentation, binary change detection, and semantic change detection, form the cornerstone of many Earth observation applications. While often treated as distinct problems, they share fundamental objectives and challenges.
Semantic segmentation (SS) in remote sensing aims to assign a thematic label to each pixel in an image, thereby partitioning it into meaningful regions corresponding to different land cover or land use classes [11]. Early approaches relied on traditional machine learning algorithms such as Support Vector Machines (SVMs) and Random Forests (RFs) operating on handcrafted features [12]. The advent of deep learning, particularly Fully Convolutional Networks (FCNs) [13], revolutionized the field, enabling end-to-end feature learning and significantly improving segmentation accuracy. Architectures like U-Net [14], with its encoder-decoder structure and skip connections, became foundational, inspiring numerous variants such as UNet++ [15] and the DeepLab series [15, 16], which introduced atrous convolutions and spatial pyramid pooling to handle scale variations.
Binary change detection (BCD) focuses on identifying areas of significant difference between two co-registered remote sensing images acquired at different times [17]. Traditional BCD methods include image differencing, principal component analysis (PCA)-based techniques [18], and change vector analysis (CVA) [19]. Deep learning has brought substantial advancements, with Siamese networks being a popular choice for learning discriminative features from image pairs [20, 21]. These networks typically process the bi-temporal images through shared-weight encoders and then compare the resulting feature maps to identify changes. U-Net-like architectures have also been adapted for BCD, often by modifying the input layers to accept concatenated bi-temporal images or difference images, and training the network to output a binary change map [22].
Semantic change detection (SCD), also known as ”from-to” change detection, provides a more detailed analysis by not only identifying where changes have occurred but also classifying the type of land cover transition (e.g., forest to urban) [23]. This task is inherently more complex than BCD. A common traditional approach is post-classification comparison (PCC), where SS is performed independently on multi-temporal images, and the resulting maps are compared [24]. However, PCC suffers from error accumulation from the individual classification steps. More integrated methods aim to perform simultaneous segmentation and change identification. Deep learning models for SCD often employ dual-branch or multi-task learning frameworks, where separate decoders might be used for segmentation at each time step and for the final change map [25].
Critically analyzing these three tasks reveals profound interrelations and shared challenges. SCD inherently builds upon the principles of SS, as identifying ”from-to” transitions requires accurate semantic understanding of the landscape at different time points. BCD can be viewed as a simplified abstraction of SCD, where all types of semantic changes are aggregated into a binary ”change” or ”no change” outcome. Thus, accurate SS is a prerequisite for high-quality SCD, and insights from BCD can potentially guide the focus of SCD algorithms. The complementary nature of these tasks is evident: SS provides the static land cover context, BCD highlights dynamic regions, and SCD offers a comprehensive narrative of landscape transformation. Information from one task can, in principle, regularize or enhance the performance of others; for instance, knowing the semantic class of a pixel can refine change boundaries, or detecting a change can prompt a more detailed semantic analysis. Despite these intrinsic relationships and the potential for synergistic modeling, there is currently a significant lack of models or frameworks that can effectively unify semantic segmentation, binary change detection, and semantic change detection within a single, cohesive architecture.
II-B Data Dimensionality in Remote Sensing Dense Prediction
Remote sensing datasets utilized for dense prediction tasks are characterized by extensive variability across temporal, spectral, and spatial dimensions. This heterogeneity stems from the diverse array of satellite and airborne sensors, each with unique acquisition parameters, mission objectives, and coverage patterns, tailored for different application scenarios and geographical regions [26]. Such variability presents a formidable challenge for developing universally applicable and robust dense prediction models.
The temporal dimension in remote sensing data exhibits significant disparities. Revisit frequency, a critical factor for monitoring dynamic phenomena, ranges from multiple observations per day with sensors like MODIS [27], to several days (e.g., Sentinel-1 and Sentinel-2, with 5-12 day repeat cycles depending on latitude and constellation status [28, 29]), to 16 days for Landsat missions [30]. Consequently, the temporal length of image sequences available for analysis can vary from bi-temporal pairs, commonly used in BCD and SCD [[20], [23]], to dense time series comprising hundreds of observations, which are invaluable for applications like agricultural monitoring [31]. Models trained on data with a specific temporal frequency or sequence length may not generalize well to datasets with different temporal characteristics without substantial retraining or adaptation.
Spectral dimensionality is another source of major variation. The number of spectral bands can range from a single panchromatic band to a few multispectral bands (e.g., 4-bands in NAIP imagery, 13 bands in Sentinel-2 MSI [29]), to hundreds of narrow, contiguous bands in hyperspectral sensors like AVIRIS [32] or the upcoming EnMAP mission [33]. Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths. This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures. However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [34].
Spatial characteristics, primarily ground sampling distance (GSD) or resolution, also vary widely. Data can range from very high resolution (VHR), with GSDs of less than 1 meter, to high resolution (HR, 1-10 meters, e.g., Sentinel-2, PlanetScope), to lower resolution. VHR imagery is suitable for detailed urban mapping or object detection [35], while medium to low-resolution data is often used for regional or global land cover mapping and monitoring [36]. The spatial extent of individual scenes also differs, impacting the scale of analysis. Models trained on imagery of a particular GSD often struggle when applied to data with significantly different spatial resolutions, as object appearance and contextual cues change drastically.
The inherent heterogeneity in temporal frequency, spectral composition, and spatial resolution across remote sensing datasets thus poses a substantial hurdle for developing universally applicable dense prediction models. Consequently, there is a pressing research gap concerning the development of adaptive model architectures or unified data processing strategies that can effectively ingest and interpret such diverse time-spectrum-space data formats at the input level for comprehensive and robust dense prediction.
III Methodology
III-A Problem Formulation
Dense prediction tasks in remote sensing generally refer to performing structured semantic inference on input images at the pixel level. While these tasks vary in nature, their core objective remains consistent: to extract rich semantic information from remote sensing data that includes temporal, spectral, and spatial dimensions, and to output structured prediction results. This paper addresses three typical tasks—semantic segmentation, semantic change detection, and binary change detection—and formulates them as a unified tensor mapping problem. Let the input remote sensing image be represented as a four-dimensional tensor $X\in\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}$, where $T_{1}$ denotes the length of the time sequence, $C_{1}$ represents the number of spectral channels, and $H_{1}$ and $W_{1}$ denote the height and width of the image, respectively. The model output is represented as a tensor $Y\in\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}}$, where $T_{2}$ represents the output time dimension, $C_{2}$ denotes the number of semantic categories, and $H_{2}$ and $W_{2}$ represent the spatial dimensions of the output image. In most cases, the input and output are spatially aligned, i.e., $H_{1}=H_{2}$ and $W_{1}=W_{2}$.
Based on this representation, the three tasks considered in this paper can be uniformly described as follows:
  • Semantic Segmentation: The model performs pixel-level multi-class classification on a remote sensing image at a given time point. This corresponds to the setup where $T_{1}=1$, $T_{2}=1$, and $C_{2}\geq 2$.
  • Semantic Change Detection: The model predicts semantic labels for each time point to analyze semantic changes between any two time points. This corresponds to $T_{2}=T_{1}$, with $C_{2}\geq 2$.
  • Binary Change Detection: The model detects changes between consecutive time points without distinguishing specific categories of change. This corresponds to $T_{2}=T_{1}-1$ and $C_{2}=2$.
It is important to note that the input data may differ significantly in terms of time length $T_{1}$, spectral channels $C_{1}$, and spatial scale $(H_{1},W_{1})$, depending on the task type or dataset. To address this variability, we propose a unified modeling framework that aims to learn a mapping function $f_{\theta}$, such that
$f_{\theta}:\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}\rightarrow\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}},$ (1)
which can adapt to any input configuration and output dense prediction results that meet the requirements of the target task. This abstract modeling provides the mathematical foundation for the modular design of the subsequent methods.
III-B Overview of the Temporal-Spectral-Spatial Unified Network (TSSUN)
The proposed Temporal-Spectral-Spatial Unified Network (TSSUN) is designed as a unified modeling framework capable of addressing diverse dense prediction tasks across varying data modalities and spatial resolutions, as illustrated in Figure 2. TSSUN employs the Temporal-Spectral-Spatial Unified Strategy (TSSUS), which systematically integrates input alignment, feature extraction, and output decoding into a cohesive processing pipeline.
In the input stage, TSSUN introduces the Spectral-Spatial Unified Module (SSUM), which encodes heterogeneous spectral and spatial data into a unified representation. By leveraging spectral and spatial priors through a guided alignment mechanism, SSUM ensures data consistency across spectral and spatial dimensions, establishing a coherent input foundation for subsequent feature modeling. During the feature extraction stage, TSSUN employs the Local-Global Window Attention (LGWA) mechanism as the primary feature modeling unit, effectively capturing multi-scale features in TSS inputs. LGWA integrates three distinct window-based attention blocks, each tailored to capture local and global features at different scales. This structure balances computational efficiency and representational power, facilitating the extraction of spatial structures across multiple scales. At the encoder-decoder junction, TSSUN incorporates the Temporal Unification Module (TUM), which focuses on temporal feature fusion. TUM mitigates temporal redundancy and adjusts the temporal length of the output based on task-specific requirements, ensuring temporal coherence in the generated predictions. In the output stage, TSSUN reintroduces SSUM in a mirrored configuration to further reinforce spectral-spatial consistency during the decoding process. By progressively decoding spectral and spatial dimensions, SSUM enables unified modeling for both semantic segmentation with arbitrary class numbers and multi-resolution spatial reconstruction, thereby enhancing TSSUN’s capability to handle multi-task, multi-modal data effectively.
Overall, TSSUN leverages TSSUS to comprehensively integrate the input, feature extraction, and output stages, establishing a unified framework for heterogeneous data representation and dense prediction. The LGWA module further enhances the model’s capacity to effectively capture both local and global contextual information, ensuring a balanced representation of multi-scale and multi-modal features. The following sections provide a detailed analysis of the two core components, TSSUS and LGWA, including their structural design and functional mechanisms.
Figure 2: Overview of the TSSUN architecture, comprising the Spectral-Spatial Unified Module (SSUM) for input alignment, Local-Global Window Attention (LGWA) blocks for feature extraction, and the Temporal Unified Module (TUM) for temporal fusion and output adaptation. The top right panel details the Dimension Unified Module (DUM), while the bottom right panel illustrates the LGWA block structure.
III-C Temporal-Spectral-Spatial Unified Strategy (TSSUS)
The Temporal-Spectral-Spatial Unified Strategy (TSSUS) is designed to decouple and unify the representation of inputs and outputs across temporal, spectral, and spatial dimensions. In the input stage, TSSUS leverages the inherent characteristics of spectral and spatial dimensions to achieve unified encoding of remote sensing imagery. Specifically, in the spectral dimension, TSSUS captures the continuity among spectral bands and incorporates spectral metadata to ensure spectral alignment. In the spatial dimension, TSSUS extracts spatial scale representations by modeling the similarity among neighboring regions and incorporating spatial metadata. By retaining the temporal dimension at the input stage, TSSUS enables independent feature extraction for remote sensing images across different time instances. At the encoder-decoder interface, TSSUS performs feature-level fusion along the temporal axis, effectively suppressing irrelevant information while adaptively adjusting the output temporal length to accommodate dense prediction tasks. In the output stage, TSSUS further leverages the structured characteristics of dense prediction tasks in remote sensing across spectral and spatial dimensions. In the spectral dimension, it models inter-class relationships to facilitate predictions with arbitrary class numbers. In the spatial dimension, it enhances the model’s generalization capability for multi-resolution data outputs by leveraging the spatial structure consistency among neighboring regions.
To facilitate the mapping from the original data space to the unified feature space across different variables (temporal, spectral, and spatial), TSSUS introduces the Dimension Unified Module (DUM), which serves as the foundational implementation for both the SSUM and the TUM. DUM employs metadata from TSS variable subsets to generate adaptive weights and biases for linear layers, thereby enabling adaptive feature mapping. As illustrated in the top-right subfigure of Figure 2, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, where $D_{1}$ varies across different tasks and scenarios, DUM maps the input tensor to a unified feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process. The detailed procedure is as follows:
  • Metadata Embedding: Metadata $M$ is subjected to positional encoding and tokenization. A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
  • Cross-Variable Relationship Modeling: The token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
  • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$. The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels. The input $X$ is first reshaped from $(D_{1},H,W)$ to $(L,D_{1})$, where $L=H\times W$, and then processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$.
The DUM structure not only adaptively maps any TSS variable subset to a unified feature in terms of shape, but also effectively preserves the relationships between remote sensing data in original space, thereby enhancing the model’s generalization capability across heterogeneous data.
III-D Local-Global Window Attention (LGWA)
Transformer networks have become a prevalent structure in remote sensing tasks due to their exceptional capacity for modeling long-range dependencies. However, the computational complexity of traditional transformers increases quadratically with sequence length, posing significant challenges for the efficient processing of large-scale remote sensing data. To address this issue, this study proposes the Local-Global Window Attention (LGWA) module, which balances computational efficiency and modeling capability by capturing fine-grained features within local windows while simultaneously incorporating global contextual information.
LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 3. These window shapes not only determine the feature sensitivity within each window but also directly impact computational efficiency. Therefore, selecting appropriate window configurations in different network blocks allows for an optimal trade-off between modeling effectiveness and computational resources. For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (2)
where $W^{Q},W^{K}\in\mathbb{R}^{d_{M}\times d_{k}}$ and $W^{V}\in\mathbb{R}^{d_{M}\times d_{v}}$ are learnable weight matrices. The attention calculation within each window focuses on extracting fine-grained features, while different window configurations provide sensitivity to varying scales.
Figure 3: Architecture of the proposed Local-Global Window Attention (LGWA)-based Transformer block, employing multiple window shapes to enhance feature extraction.
Next, attention scores are computed using the scaled dot-product attention mechanism:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (3)
To further enhance the feature representation, the LGWA module employs the multi-head attention strategy, allowing each head to independently execute the above process and then concatenate the results:
$\text{MultiHead}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h})W^{O},$ (4)
where $H_{i}=\text{Attn}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$, and $W^{O}\in\mathbb{R}^{hd_{v}\times d_{M}}$ projects the concatenated output back to the original dimension.
By adopting this modular design, the LGWA module effectively captures fine-grained features within local windows while integrating global contextual information. The flexible configuration of window shapes further optimizes computational efficiency, enabling the module to handle multi-scale remote sensing scenarios effectively.
IV Experiments and Results
To validate the adaptability of TSSUN to arbitrary Temporal-Spectral-Spatial inputs and its capability to concurrently perform semantic segmentation, semantic change detection, and binary change detection tasks with support for a variable number of output classes, we conducted experiments on a total of six datasets across building and Land Cover/Land Classification (LULC) scenarios. For each scenario, a unified TSSUN model was trained using the combined training sets from all datasets within that scenario, and its performance was evaluated on their respective test sets.
In the building scenario, we selected the WHU, WHU-CD, LEVIR-CD, and TSCD datasets. While these datasets share the same number of channels in their input and label data, they exhibit variations in temporal and spatial dimensions, leading to differences in dense prediction task types. Specifically, the WHU dataset corresponds to a single-temporal building semantic segmentation task. The WHU-CD dataset is associated with a bi-temporal building semantic change detection task. The LEVIR-CD dataset is used for a bi-temporal building binary change detection task. The TSCD dataset pertains to a multi-temporal building binary change detection task, as summarized in Table I.
In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets. These datasets differ in their temporal, spectral dimensions for both input and label data. The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task, while the Dynamic EarthNet dataset is used for multi-temporal semantic change detection, semantic segmentation and binary change detection tasks, as detailed in Table I.
The experiments across multiple datasets in the building scenario were primarily designed to verify TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spatial dimensions within this context. Furthermore, these experiments aimed to confirm its ability to simultaneously handle semantic segmentation, semantic change detection, and binary change detection tasks. The datasets in the LULC scenario were mainly used to validate TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spectral dimensions. Additionally, these experiments were intended to demonstrate its capacity to concurrently manage semantic segmentation, binary change detection and semantic change detection tasks while supporting a flexible number of output classes.
TABLE I: Brief Introduction of The Experimental Datasets.
Name Scene Task Input T Output T Input C Output C Image Size Resolution Images
WHU Building SS 1 1 3 1 512×512 0.3 8189
WHU-CD Building SCD 2 2 3 1 32207×15354 0.075 1
LEVIR-CD Building BCD 2 1 3 1 1024×1024 0.5 445
TSCD Building BCD 4 3 3 1 256×256 0.5 2700
LoveDA Urban LULC SS 1 1 3 7 1024×1024 0.3 5987
DynamicEarthnet LULC SCD&SS&BCD 24 24 4 6 1024×1024 3 54750
IV-A Datasets
We offer a brief description of the experimental building and LULC scene datasets in Table I.
1) Building scene datasets: The WHU Building dataset [37] is divided into two main components: one containing satellite imagery and another composed of aerial photos. In our study, we utilize the aerial photo subset, which consists of 8,189 images. These images are split into 4,736 for training, 1,036 for validation, and 2,416 for testing, each with a spatial resolution of 0.3 meters. In total, this subset represents over 22,000 buildings covering an area in excess of 450 square kilometers. Our experiments were conducted using the original partitioning scheme and image dimensions (512×512) as specified by the WHU dataset.
The WHU-CD dataset [37] includes bitemporal very high-resolution (VHR) aerial images taken in 2012 and 2016, which clearly highlight major changes in building structures. The dataset is partitioned into non-overlapping patches of 1024×1024 pixels. These patches are further allocated into training, validation, and test sets following a 7:1:2 ratio.
The LEVIR-CD dataset [25] is an extensive resource for change detection, comprising VHR Google Earth images with a resolution of 0.5 m/pixel. These images capture a variety of building transformations over periods ranging from 5 to 14 years, with a particular emphasis on construction and demolition events. The bitemporal images have been expertly annotated using binary masks, where a label of 1 denotes a change and 0 signifies no change. In total, there are 31,333 labeled instances of building modifications. Our experimental setup used the dataset’s original image dimensions of 1024×1024 and adhered to the provided data partitioning scheme.
The TSCD dataset [38] is constructed from WorldView-2 satellite imagery with a spatial resolution of approximately 0.5 m/pixel, acquired in 2016, 2018, 2020, and 2022. To mitigate external influences, the images underwent co-registration using manually selected control points and resampling to ensure a consistent coordinate framework. Building footprints were densely labeled for each temporal phase. Subsequently, three sets of change labels (2016–2018, 2018–2020, 2020–2022) were generated by performing differential operations on adjacent building distribution maps. The final TSCD dataset was created through uniform cropping and partitioning of these original images and derived labels.
2) LULC scene datasets: The LoveDA dataset [39] consists of 5,987 high-resolution optical remote sensing images (with a ground sampling distance of 0.3 m) each sized at 1024×1024 pixels. It covers seven land cover classes: building, road, water, barren, forest, agriculture, and background. The dataset is divided into 2,522 training images, 1,669 images for validation, and 1,796 images for testing, all drawn from two distinct scenes—urban and rural—from three Chinese cities: Nanjing, Changzhou, and Wuhan. The dataset poses considerable challenges due to the presence of multiscale objects, complex backgrounds, and uneven class distribution.
The DynamicEarthnet dataset [40] comprises 55 daily Sentinel-2 Image Time Series (SITS) collected globally between January 1, 2018, and December 31, 2019. For each month, data from the first day is annotated, which results in 24 ground truth segmentation maps per Area of Interest (AoI). Each image is 1024×1024 pixels and multi-spectral, containing four channels (RGB plus near-infrared). The annotations cover general land-use and land-cover categories: impervious surface, agriculture, forest, wetlands, soil, and water. The ’snow’ class appears in only a few AoIs and has been excluded from this study.
IV-B Baseline
To evaluate the effectiveness of the proposed TSSUN, we conducted comparative experiments with various benchmark methods on the building and LULC scene datasets. The benchmark methods tested on the same dataset are based on the same splitting of the dataset and use the same data.
On the four building scene datasets, the compared CNN-based models include FCN [13], SegNet [41], U-Net [14], PSPNet [42], HRNet [43], MA-FCN [44], Deeplabv3+ [16], ResUNet [45], MAPNet [46], D-LinkNet [47], SIINet [48], FC-EF [20], FC-Siam-Diff [20], FC-Siam-Conc [20], STANet [25], DTCDSCN [49], SNUNet [50], CDNet [51], DDCNN [52], DASNet [53], DSIFN [54], HANet [55], USSFCNet [56], and SEIFNet [57], the compared transformer-based models include Segformer [58], ChangeFormer [59] and A2Net [60], and the CNN-transformer hybrid models include BDTNet [61], TransUNet [62], CMTFNet [63], BIT [64], MTCNet [65], MSCANet [66], AMTNet-50 [67], Contrast-COUD [38] and TS-COUD [38].
On the two LULC scene datasets, the compared CNN-based models include U-Net [14], Deepabv3+ [16], DANet [68], ResUNet-a [45], DASSN [69], HCANet [70], RAANet [71], A2-FPN [72] and CAC [73], the compared transformer-based models include LANet [74], SCAttNet [75], and TSViT [76], the CNN-transformer hybrid models include SAPNet [77], SSCBNet [78], UTAE [79], A2Net [80], SCanNet [81] and TSSCD [82].
IV-C Implementation Details
1) Data Augmentation: To validate the proposed methods, we adopted a minimalistic yet effective data augmentation strategy, deliberately refraining from complex augmentation schemes. Specifically, the employed transformations were limited to horizontal/vertical flipping (probability = 0.5) and transposition (probability = 0.5).
2) Training and Inference: The TSSUN model was implemented using PyTorch [83] and executed on a single RTX A100 GPU (80G). Due to the heterogeneous image resolutions across the datasets, the batch size was set to 16 for the four building scene datasets and 4 for the two LULC scene datasets. Our optimization strategy combined binary cross-entropy loss with Dice coefficient loss, facilitating a balanced performance optimization. The AdamW optimizer [84] was initialized with a learning rate of 0.0001 and a weight decay of 0.001. A learning rate scheduler was employed to reduce the learning rate by a factor of 0.1 if no increase in the mean F1-score was observed on the aggregate validation set for 5 consecutive epochs. The training process spanned 100 epochs, ensuring robust convergence, and the best performing checkpoints—corresponding to the maximum mean F1-scores achieved—were retained for the testing phase. Furthermore, in order to ensure comparability with existing methodologies, all models were initialized using the default PyTorch settings across all datasets.
3) Evaluation Metrics: The performance of the proposed models was quantitatively assessed using five principal metrics: overall accuracy (OA), precision (P), recall (R), F1-score, and intersection over union (IoU). For multi-temporal tasks and multi-category tasks, the average F1-score (AF) and mean IoU (mIoU) will be used. In addition, following the settings of the DynamicEarthnet dataset [40], we use the semantic change segmentation (SCS) metric, classagnostic binary change score (BC) and semantic segmentation score among changed pixels (SC) metrics to evaluate the model’s performance on this dataset. Precision provides a measure of the incidence of false positives, while recall reflects the occurrence of false negatives. Due to the inherently inverse relationship between these two metrics, simultaneously securing high scores for both represents a significant challenge. The F1-score, calculated as the harmonic mean of precision and recall, offers a balanced assessment. The IoU metric evaluates spatial accuracy by quantifying the overlap between the predicted and ground truth changed pixels relative to their union.
IV-D Ablation Study
To ascertain the efficacy of the proposed Temporal-Spectral-Spatial Unified Strategy (TSSUS) and Local-Global Window Attention (LGWA), ablation studies were performed on the TSCD dataset. As comparative baselines, TSSUS was replaced with a strategy involving the direct unification of temporal, spectral, and spatial dimensions at both the input and output stages ( referred to as ”direct unification”). Concurrently, LGWA was substituted with standard global attention. This experimental design serves to highlight the distinct advantages of TSSUS in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
TABLE II: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
Strategy Attention F1 IoU OA
TSSUS LGWA 66.48 49.79 98.05
Direct Unification LGWA 63.22 46.22 94.72
TSSUS Global Attention 65.91 49.15 97.21
Direct Unification Global Attention 62.87 45.85 94.38
Table 2 presents the results of these ablation studies, indicating that both TSSUS and LGWA yield superior performance compared to their respective baseline alternatives. Specifically, TSSUS distinguishes itself from the direct unification approach. It begins by unifying the spectral and spatial dimensions at the input stage. Then, within the encoder, it independently extracts features for each temporal instance (i.e., each remote sensing image). Subsequently, these temporally distinct features are integrated via feature-level fusion at the interface between the encoder and decoder. This methodology—processing temporal information independently before fusion—effectively minimizes interference from irrelevant pixel information within the remote sensing imagery. Furthermore, TSSUS dynamically adjusts the temporal length of feature fusion to align with the specific requirements of the target dense prediction task. Following this, it optimizes the output for each time step independently within the decoder, thereby significantly enhancing the overall model efficacy.
In a similar vein, LGWA offers advantages over standard global self-attention. By employing a combination of variously shaped local windows alongside global self-attention mechanisms, LGWA is capable of concurrently extracting a rich tapestry of local features and comprehensive global context. This simultaneous extraction of multi-level features is crucial for enabling the model to effectively perform dense prediction tasks across various scales.
IV-E Overall Comparison
IV-E1 Building
The efficacy of the proposed Temporal-Spectral-Spatial Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection. Our method was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed TSSUN achieves state-of-the-art performance. As detailed in Table III, TSSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%. This superior performance can be attributed to TSSUN’s Temporal-Spectral-Spatial Unified Strategy, particularly its spatial unification component, which effectively processes varying spatial resolutions, and the Local-Global Window Attention mechanism. The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
TABLE III: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
Methods P (%) R (%) F1 (%) IoU (%)
FCN [13] 92.29 92.84 92.56 86.16
SegNet [41] 93.42 91.71 92.56 86.15
U-Net [14] 94.50 90.88 92.65 86.31
PSPNet [42] 93.19 94.21 93.70 88.14
HRNet [43] 91.69 92.85 92.27 85.64
MA-FCN [44] 94.75 94.92 94.83 90.18
Deeplabv3+ [16] 94.31 94.53 94.42 89.43
ResUNet [45] 94.49 94.71 94.60 89.75
MAP-Net [46] 93.99 94.82 94.40 89.40
Segformer [58] 94.72 94.42 94.57 89.70
TransUNet [62] 94.05 93.07 93.56 87.89
CMTFNet [63] 90.12 95.21 92.59 86.21
TSSUN 95.71 94.87 95.29 91.00
For the bi-temporal semantic change detection task on the WHU-CD dataset, TSSUN demonstrates leading performance against other SOTA methods, as shown in Table IV. The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points. TSSUN excels here due to its inherent design for unified temporal and spectral modeling. The TSSUS component allows the network to effectively learn temporal evolutionary patterns and relationships between different land cover classes, while the LGWA mechanism enhances the discrimination of subtle yet significant semantic changes from unchanged areas, leading to more precise change maps.
TABLE IV: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
Methods P (%) R (%) F1 (%) IoU (%)
FCN [13] 79.35 77.82 78.58 64.71
SegNet [41] 85.20 86.21 85.70 74.98
Deeplabv3+ [16] 89.24 90.91 90.07 81.93
U-Net [14] 83.19 84.02 83.60 71.83
PSPNet [42] 84.85 82.09 83.45 71.60
HRNet [43] 86.77 85.92 86.34 75.97
MA-FCN [44] 86.10 89.92 87.97 78.52
Segformer [58] 90.45 88.93 89.68 81.30
TransUNet [62] 93.82 89.33 91.52 84.37
RSM-CD 93.07 91.20 92.13 85.40
In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, TSSUN surpasses existing methods. Table V shows that our method achieves the highest F1-score of 91.59% and an IoU of 84.49%. The strength of TSSUN on this dataset lies in its robust temporal modeling capabilities, facilitated by TSSUS, which allows for consistent feature representation across different time points. Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
TABLE V: Accuracy Comparison on the LEVIR-CD Dataset. The Best Values Are Highlighted in Bold.
Methods P (%) R (%) F1 (%) IoU (%)
FC-EF [20] 86.91 80.17 83.40 71.53
FC-Siam-Diff [20] 89.53 83.31 86.31 75.91
FC-Siam-Conc [20] 91.99 76.77 83.69 71.96
DTCDSCN [49] 88.53 86.83 87.67 78.05
DSIFN [54] 94.02 82.93 88.13 78.77
STANet [25] 83.81 91.00 87.26 77.39
SNUNet [50] 89.18 87.17 88.16 78.83
HANet [55] 91.21 89.36 90.28 82.27
CDNet [51] 91.60 86.50 88.98 80.14
DDCNN [52] 91.85 88.69 90.24 82.22
BIT [64] 89.24 89.37 89.30 80.68
ChangeFormer [59] 92.05 88.80 90.40 82.47
MTCNet [65] 90.87 89.62 90.24 82.22
MSCANet [66] 91.30 88.56 89.91 81.66
AMTNet-50 [67] 91.82 89.71 90.76 83.08
TSSUN 93.17 90.07 91.59 84.49
Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, TSSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%. The TSCD dataset requires robust modeling of temporal dependencies across multiple observations. TSSUS is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of land cover. This, combined with the LGWA’s capacity to aggregate contextual information effectively over extended temporal sequences, enables TSSUN to accurately identify changes in complex, evolving landscapes where other methods might falter due to the increased temporal dimensionality.
TABLE VI: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
Methods F1 (%) IoU (%) OA (%)
FC-EF [20] 53.39 37.86 97.03
FC-Siam-Conc [20] 41.51 28.05 96.60
FC-Siam-Diff [20] 39.65 26.83 96.48
SNUNet-CD [50] 63.22 47.22 97.61
USSFCNet [56] 55.68 39.80 97.30
A2Net [60] 53.16 37.19 97.14
SEIFNet [57] 60.37 44.01 97.41
Contrast-COUD [38] 64.35 48.45 97.72
TS-COUD [38] 65.24 49.33 97.90
TSSUN 66.48 49.79 98.05
Figure 4: Sample inference results on for building scene datasets. The input images, ground truths and predictions are shown in the first, second and third rows, respectively. Red areas denote false positives and blue areas denote false negatives. (a) WHU dataset sample. (b) WHU-CD dataset sample. (c) LEIVR-CD dataset sample. (d) TSCD dataset sample.
Figure 4 presents a qualitative comparison of our method against other SOTA approaches on representative samples from all four datasets. Visually, TSSUN consistently produces more accurate and complete segmentation and change maps. The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations. For instance, in building segmentation on the WHU dataset (Figure 4(a)), TSSUN generates sharper edges and more complete building shapes. Similarly, for change detection tasks on WHU-CD (Figure 4(b)), LEVIR-CD (Figure 4(c)), and TSCD (Figure 4(d)), our method demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas. This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of TSSUS and LGWA, to learn highly discriminative features and effectively model contextual relationships across the temporal, spectral, and spatial dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
IV-E2 LULC
The efficacy of the proposed Temporal-Spectral-Spatial Unified Network is rigorously evaluated against SOTA methods on two challenging benchmark datasets: LoveDA for single-temporal semantic segmentation and DynamicEarthNet for multi-temporal land cover classification.
On the LoveDA dataset, TSSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII. This leading performance can be attributed to TSSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information. This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery. Furthermore, the Temporal-Spectral-Spatial Unified Strategy enables robust feature representation by adeptly unifying spatial and spectral information, which is vital for handling the heterogeneity present in such diverse urban and rural landscapes with uneven class distributions.
For the multi-temporal DynamicEarthNet dataset, TSSUN again surpasses existing methods, yielding the top scores across all reported metrics: Semantic Change Segmentation (SCS) score of 29.9, class-agnostic Binary Change score (BC) of 38.9, and mIoU of 54.7 as shown in Table VIII. The success of TSSUN on this dataset underscores the effectiveness of the TSSUS in explicitly modeling the temporal dimension. By leveraging temporal meta-information to ensure consistency across sequences of varying lengths, TSSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet. This robust temporal modeling, distinct from its spatial feature extraction strengths highlighted in the LoveDA analysis, allows for consistent and accurate classification and change analysis across the 24 annotated time steps.
TABLE VII: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
Methods F1-score per category (%) AF OA mIoU
Background Building Road Water Barren Forest Agriculture
U-Net [14] 50.21 54.74 56.38 77.12 18.09 48.93 66.05 53.07 51.81 47.84
DeepLabV3+ [16] 52.29 54.99 57.16 77.96 16.11 48.18 67.79 53.50 52.30 47.62
DANet [68] 54.47 61.02 63.37 79.17 26.63 52.28 70.02 58.14 54.64 50.18
ResUNet-a [45] 59.16 64.08 66.73 81.01 32.23 55.81 75.79 62.12 59.65 54.16
DASSN [69] 57.95 66.90 68.63 76.64 44.35 54.96 70.49 62.85 60.35 55.42
HCANet [70] 66.39 70.76 75.11 88.29 51.14 63.92 81.07 70.95 69.47 62.77
RAANet [71] 55.02 62.19 65.58 81.03 29.25 54.11 74.07 60.18 58.95 53.93
SCAttNet [75] 65.95 71.88 77.04 86.61 50.79 61.19 82.00 70.78 67.31 61.09
A2FPN [72] 65.17 73.32 75.19 88.01 48.82 59.96 79.71 70.03 66.89 61.14
LANet [74] 67.04 74.19 77.54 87.54 52.23 64.78 80.80 72.02 69.11 62.16
MSAFNet [85] 65.51 73.71 75.59 88.47 49.08 60.28 80.13 70.40 67.17 60.76
CLCFormer [86] 67.17 74.34 77.69 87.71 52.34 64.91 80.96 72.16 69.37 63.85
SAPNet [77] 67.50 75.06 78.12 88.35 53.10 65.50 81.30 73.04 70.12 63.45
SSCBNet [78] 68.25 75.90 79.00 89.10 54.00 66.20 82.15 74.05 70.95 64.58
TSSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
TABLE VIII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.
[…]
CAC [73] 17.7 10.7 24.7 37.9
U-Net [14] 17.3 10.1 24.4 37.6
TSViT [76] 23.0 34.1 11.8 50.5
UTAE [79] 25.9 38.0 13.8 53.7
A2Net [80] 22.2 32.9 11.5 47.2
SCanNet [81] 24.8 35.8 13.9 53.0
TSSCD [82] 12.0 19.4 4.7 33.9
TSSUN 29.9 38.9 21.8 54.7
V DISCUSSION
The proliferation of diverse remote sensing data sources necessitates models capable of unified processing across varied data characteristics and task requirements. This research confronts the critical challenge of existing deep learning models in remote sensing dense prediction, which are typically constrained to fixed input-output configurations and struggle to generalize across the inherent heterogeneity in temporal length, spectral channels, and spatial resolution of imagery, as well as diverse task specifications such as semantic segmentation, semantic change detection, and binary change detection. This rigidity imposes significant burdens in terms of retraining and fine-tuning for each specific scenario, limiting practical applicability in dynamic operational environments.
To surmount this limitation, we proposed the Temporal-Spectral-Spatial Unified Network, a novel architecture engineered for comprehensive unification. The cornerstone of TSSUN is the Temporal-Spectral-Spatial Unified Strategy, which systematically decouples and standardizes data representations. On the input side, TSSUS leverages dimension-specific meta-information to harmonize temporal sequences of varying lengths, align spectral inputs with differing channel counts, and reconcile features from disparate spatial resolutions. This is achieved by modeling the continuity in temporal evolution, the inherent relationships between spectral bands, and the local similarity in spatial regions. Concurrently, on the output side, TSSUS exploits shared structural elements among the three primary dense prediction tasks. It supports variable temporal output lengths by identifying common temporal patterns, accommodates an arbitrary number of output classes by capturing inter-class relationships, and ensures spatial consistency across resolutions. Furthermore, the integration of a Local-Global Window Attention mechanism enhances TSSUN’s capacity. LGWA efficiently processes local contextual information through overlapping windowed attention blocks and aggregates global context via a subsequent global attention layer, striking a balance between computational load and feature representation power essential for complex remote sensing scenes.
The efficacy of the TSSUN framework in addressing the identified research problem is substantiated by extensive experimental validation. Results across multiple datasets, encompassing diverse Temporal-Spectral-Spatial configurations and tasks like building detection and land use/land cover classification, demonstrate that a single, pre-trained TSSUN model can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks, consistently achieving performance comparable to or exceeding state-of-the-art specialized models.
Despite these advancements, the current TSSUN framework possesses certain limitations. Firstly, the TSSUS component relies on the availability and accuracy of explicit meta-information (e.g., acquisition dates, sensor band specifications, ground sampling distance) to guide the unification across temporal, spectral, and spatial dimensions. The performance may be suboptimal if such meta-information is imprecise, incomplete, or unavailable, potentially hindering the model’s adaptability in data-scarce or poorly documented scenarios. Secondly, while the LGWA mechanism is designed for efficiency, the computational demand of its global attention component, although optimized, might still present a scalability challenge when processing extremely large-scale images or exceptionally long and dense temporal sequences, potentially requiring substantial memory and processing resources.
Future research will focus on mitigating these limitations and expanding the capabilities of TSSUN. One promising direction involves developing methods for the implicit learning or unsupervised inference of the requisite meta-information directly from the remote sensing data, thereby reducing dependency on external inputs and enhancing robustness. Another avenue for exploration is the refinement of the LGWA module, possibly by incorporating more sophisticated hierarchical attention structures or adaptive computational pruning techniques to further improve its efficiency and scalability for massive datasets. Additionally, future work could extend the unification paradigm of TSSUN to encompass a broader array of remote sensing tasks and potentially integrate other data modalities, such as SAR or LiDAR, within a unified analytical framework.
VI CONCLUSION
This paper addressed the critical challenge of processing heterogeneous remote sensing data for diverse dense prediction tasks, a common hurdle due to variations in temporal, spectral, and spatial characteristics. We introduced the Temporal-Spectral-Spatial Unified Network, a novel deep learning architecture. TSSUN demonstrates remarkable flexibility by handling arbitrary input configurations across these three dimensions and unifying key dense prediction tasks: semantic segmentation, semantic change detection, and binary change detection, with adaptable output class settings. The proposed Temporal-Spectral-Spatial Unified Strategy enables this unification by effectively decoupling and modeling data representations guided by meta-information. Furthermore, the designed Local-Global Window Attention mechanism significantly improves feature extraction by capturing both local and global contextual information efficiently. Comprehensive experimental validation on multiple datasets confirmed that a single TSSUN model consistently adapts to varied data inputs and task requirements, achieving or exceeding state-of-the-art performance. The TSSUN framework represents a substantial advancement towards universal remote sensing models, mitigating the need for task-specific designs and extensive retraining, thus paving the way for more scalable and cost-effective analysis of complex, multi-source geospatial data and fostering broader application of remote sensing technologies.

### v2

@@
-Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction
+Spatial-Temporal-Spectral Unified Modeling for Remote Sensing Dense Prediction
 Abstract
-The proliferation of diverse remote sensing data has spurred advancements in dense prediction tasks, yet significant challenges remain in handling data heterogeneity.
-Remote sensing imagery exhibits substantial variability across temporal, spectral, and spatial (TSS) dimensions, complicating unified data processing.
-Current deep learning models for dense prediction tasks, such as semantic segmentation and change detection, are typically tailored to specific input-output configurations.
-Consequently, variations in data dimensionality or task requirements often lead to significant performance degradation or model incompatibility, necessitating costly retraining or fine-tuning efforts for different application scenarios.
-This paper introduces the Temporal-Spectral-Spatial Unified Network (TSSUN), a novel architecture designed for unified representation and modeling of remote sensing data across diverse TSS characteristics and task types.
-TSSUN employs a Temporal-Spectral-Spatial Unified Strategy that leverages meta-information to decouple and standardize input representations from varied temporal, spectral, and spatial configurations, and similarly unifies output structures for different dense prediction tasks and class numbers.
-Furthermore, a Local-Global Window Attention mechanism is proposed to efficiently capture both local contextual details and global dependencies, enhancing the model’s adaptability and feature extraction capabilities.
-Extensive experiments on multiple datasets, encompassing tasks like building detection and land use/land cover classification with diverse TSS configurations, demonstrate that a single TSSUN model effectively adapts to heterogeneous inputs and unifies various dense prediction tasks.
-The proposed approach consistently achieves or surpasses state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications without requiring task-specific modifications.
+The proliferation of multi-source remote sensing data has propelled the development of deep learning for dense prediction, yet significant challenges in data and task unification persist.
+Current deep learning architectures for remote sensing are fundamentally rigid.
+They are engineered for fixed input-output configurations, restricting their adaptability to the heterogeneous spatial, temporal, and spectral dimensions inherent in real-world data.
+Furthermore, these models neglect the intrinsic correlations among semantic segmentation, binary change detection, and semantic change detection, necessitating the development of distinct models or task-specific decoders.
+This paradigm is also constrained to a predefined set of output semantic classes, where any change to the classes requires costly retraining.
+To overcome these limitations, we introduce the Spatial-Temporal-Spectral Unified Network (STSUN) for unified modeling.
+STSUN can adapt to input and output data with arbitrary spatial sizes, temporal lengths, and spectral bands by leveraging their metadata for a unified representation.
+Moreover, STSUN unifies disparate dense prediction tasks within a single architecture by conditioning the model on trainable task embeddings.
+Similarly, STSUN facilitates flexible prediction across any set of semantic categories by integrating trainable category embeddings as metadata.
+Extensive experiments on multiple datasets with diverse STS configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions.
+The proposed approach consistently achieves state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications.
 Index Terms:
-Remote Sensing, Dense Prediction, Unified Network, Change Detection, Semantic Segmentation, Deep learning
+Remote Sensing, Dense Prediction, Data Unification, Task Unification, Change Detection, Semantic Segmentation, Deep learning
 I INTRODUCTION
@@
 Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7].
-Remote sensing imagery exhibits high heterogeneity across three key dimensions—temporal, spectral, and spatial—posing significant challenges for unified processing due to variations in temporal length, number of spectral channels, and spatial resolution in practical applications [8].
-From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, semantic change detection, and binary change detection.
-These tasks can be formally defined as follows: given a remote sensing image of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the spatial dimensions of the input and output, respectively.
+Remote sensing imagery exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands in practical applications [8].
+Remote sensing dense prediction refers to tasks where the goal is to produce a pixel-wise prediction, assigning a specific label to every pixel in the input satellite imagery [9].
+From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection.
+These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively.
 Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$.
-In semantic segmentation, the model performs multi-class land cover extraction from a single time-point image, corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points, corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points, corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$.
-In recent years, deep learning methods, supported by the growing abundance of remote sensing data, have achieved notable success in dense prediction tasks, leading to the emergence of a wide range of high-performing models across all three task categories.
-However, existing models are typically designed for fixed input-output configurations, i.e., specific $(T_{1},C_{1},H_{1},W_{1})$ and $(T_{2},C_{2},H_{2},W_{2})$.
-Even slight variations in any of these dimensions can lead to significant performance degradation or complete incompatibility.
-In particular, changes in $T_{1}$, $C_{1}$, and $(H_{1},W_{1})$ reflect the diversity of remote sensing data along the Temporal- Spectral-Spatial (TSS) dimensions; variations in $T_{2}$ correspond to different task types, while differences in $C_{2}$ relate to varying classification requirements.
-Given that remote sensing applications often differ across these dimensions, models trained in one scenario typically struggle to generalize to others, requiring additional training or fine-tuning, which introduces considerable computational and time costs.
-Figure 1: Illustration of Temporal-Spectral-Spatial Unified Network.
-TSSUN is capable of handling input and output with arbitrary temporal, spectral, and spatial dimensions, supports dense prediction for any number of classes, and unifies semantic segmentation, binary change detection and semantic change detection tasks.
-To address these issues, this study propose a Temporal-Spectral-Spatial Unified Network (TSSUN) that enables unified representation and modeling of remote sensing data across the temporal, spectral, and spatial dimensions, as shown in Figure 1.
-TSSUN is highly flexible, capable of handling any combination of $(T_{1},C_{1},H_{1},W_{1})$ inputs and $(T_{2},C_{2},H_{2},W_{2})$ outputs.
-It supports inputs with varying temporal lengths, spectral bands, and spatial resolutions, and unifies all three dense prediction tasks.
-Moreover, it can produce outputs with an arbitrary number of semantic classes and maintains strong performance across diverse configurations.
-Specifically, we introduce the Temporal-Spectral-Spatial Unified Strategy (TSSUS) to decouple and unify the input and output representations across the TSS dimensions.
-At the input stage, TSSUS leverages inherent spectral and spatial characteristics to achieve unified encoding of remote sensing imagery.
-Along the spectral dimension, it exploits the continuity across spectral bands and integrates spectral metadata to accomplish spectral harmonization.
-In parallel, along the spatial dimension, it harnesses the similarity among nearby regions, supplemented with spatial metadata, to produce a unified representation of spatial scales.
-By preserving the time dimension at the input, TSSUS is able to independently extract features from remote sensing images across distinct temporal instances.
-At the encoder–decoder junction, TSSUS facilitates feature-level fusion along the temporal axis, effectively mitigating interference from irrelevant information while adapting the output temporal length to suit the dense prediction task at hand [9, 10].
-At the output stage, TSSUS further capitalizes on the structured nature of remote sensing dense prediction tasks in both spectral and spatial dimensions.
-In the spectral dimension, by modeling inter-class correlations, it enables prediction over an arbitrary number of classes.
-Meanwhile, in the spatial dimension, by exploiting the spatial structural consistency among neighboring regions, it significantly enhances the model’s generalization capability when outputting data at varying spatial resolutions.
-To further improve the model’s ability to capture diverse TSS combinations, we design a Local-Global Window Attention (LGWA) mechanism.
+In semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$.
+Therefore, the temporal dimension $T_{1},T_{2}$ is strongly correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is strongly correlated with the set of predicted semantic categories.
+Figure 1: Illustration of Spatial-Temporal-Spectral Unified Network.
+STSUN is capable of handling input and output with arbitrary spatial, temporal and spectral dimension configurations, unifies semantic segmentation, binary change detection and semantic change detection tasks with flexible category set.
+While deep learning methods have achieved significant progress in remote sensing dense prediction, yielding high-performing models across semantic segmentation, binary change detection, and semantic change detection tasks [12, 13, 14, 15], current architectures exhibit several limitations.
+1.
+Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H1,W1),(H2,W2))$, temporal lengths $(T1,T2)$ and spectral bands $(C_{1},C_{2})$.
+This rigidity restricts their adaptability to the diverse and heterogeneous remote sensing datasets encountered in practical applications, such as diverse satellite sensor data with varying spectral bands or multi-temporal imagery sequences with varying temporal lengths.
+2.
+Fixed Task: These models neglect the inherent correlations among semantic segmentation, binary change detection, and semantic change detection tasks.
+This oversight often mandates the development of distinct models or dedicated task-specific decoders for each task, hindering unified processing pipelines.
+3.
+Fixed Category Set: Existing approaches are generally constrained to dense prediction with a fixed set of output categories.
+Consequently, any alteration to the target dense prediction schema, even a slight change, often result in substantial performance degradation or complete incompatibility, necessitating extensive retraining or fine-tuning, which incurs considerable computational and temporal overhead.
+To overcome these limitations, this study introduces the Spatial-Temporal-Spectral Unified Network (STSUN), as illustrated in Figure 1.
+STSUN offers a novel framework for unified representation and modeling of remote sensing data across diverse spatial, temporal and spectral dimensions, solving the above issues in the following ways.
+1.
+Unified Representation: STSUN effectively addresses the variability in data characteristics by leveraging inherent metadata from each dimension to achieve a unified representation.
+This capability allows STSUN to seamlessly adapt to input and output data with arbitrary image sizes, temporal lengths, and spectral bands, thereby leveraging a wide range of data to build high-performance unified models for dense prediction.
+2.
+Unified Tasks: STSUN unifies semantic segmentation, binary change detection, and semantic change detection tasks within a single framework, which is accomplished by using a predefined and trainable task embedding set.
+By incorporating a selected task embedding and output temporal length into the output temporal dimension as metadata, the model is explicitly instructed to execute on the exact dense prediction task and the necessary output temporal length.
+This approach eliminates the need for separate models or task-specific decoders, enabling effective multi-task joint modeling and leveraging extensive data from multiple dense prediction tasks to improve model performance.
+3.
+Flexible Category Set: STSUN facilitates flexible prediction across an arbitrary set of semantic categories, which is accomplished by using a predefined and trainable category embedding set.
+The selected category subset is integrated into the output channel dimension as metadata, which guides the model to identify the specific classes needed for a given prediction.
+This enables the model to adapt to multiple subsets of prediction categories without fine-tuning, applying to diverse remote sensing scenarios with various semantic categories.
+To further improve the model’s ability to capture diverse STS combinations, we design a Local-Global Window Attention (LGWA) mechanism.
 This module efficiently extracts local features using three overlapping window-based attention blocks with different shapes, followed by a global attention block that aggregates information at the global level.
-This design achieves a balance between computational efficiency and expressive power, enabling collaborative modeling of local and global features and significantly boosting the model’s performance in complex remote sensing tasks.
-In summary, the main contributions of this work are as follows:
+This design achieves a balance between computational efficiency and expressive power, enabling collaborative modeling of local and global features and boosting the model’s performance in complex remote sensing tasks.
+The main contributions of this work are as follows:
   1. 1.
-    A Temporal-Spectral-Spatial Unified Network is proposed to enable unified modeling across temporal, spectral, and spatial dimensions for dense prediction in remote sensing.
-TSSUN supports arbitrary TSS input configurations and accommodates various tasks, including semantic segmentation, semantic change detection, and binary change detection with flexible output class settings.
+    We propose the Spatial-Temporal-Spectral Unified Network, which accommodates arbitrary input and output configurations across spatial, temporal, and spectral dimensions for the first time.
+This addresses the prevalent issue of model rigidity in handling heterogeneous remote sensing data.
   2. 2.
-    A Local-Global Window Attention mechanism is designed to efficiently capture both local and global contextual features, improving performance across a wide range of remote sensing prediction tasks.
+    We introduce a unified task execution method that leverages trainable task embeddings to perform semantic segmentation, binary change detection, and semantic change detection within a single architecture, thereby obviating the need for separate models or task-specific heads.
   3. 3.
-    Extensive experiments are conducted on multiple datasets with diverse TSS configurations, including building detection and land use/land cover (LULC) classification.
-The results show that a single TSSUN model can adapt to heterogeneous inputs, unify multiple task types, and consistently achieve or exceed state-of-the-art (SOTA) performance.
+    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for any specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
+  4. 4.
+    We develop a Local-Global Window Attention mechanism that efficiently captures both local and global contextual features.
+This design enhances the model’s feature extraction capabilities and improves performance across a wide range of remote sensing prediction tasks.
 II Related Works
-II-A Task Unification in Remote Sensing Dense Prediction
-Dense prediction tasks in remote sensing, particularly semantic segmentation, binary change detection, and semantic change detection, form the cornerstone of many Earth observation applications.
-While often treated as distinct problems, they share fundamental objectives and challenges.
-Semantic segmentation (SS) in remote sensing aims to assign a thematic label to each pixel in an image, thereby partitioning it into meaningful regions corresponding to different land cover or land use classes [11].
-Early approaches relied on traditional machine learning algorithms such as Support Vector Machines (SVMs) and Random Forests (RFs) operating on handcrafted features [12].
-The advent of deep learning, particularly Fully Convolutional Networks (FCNs) [13], revolutionized the field, enabling end-to-end feature learning and significantly improving segmentation accuracy.
-Architectures like U-Net [14], with its encoder-decoder structure and skip connections, became foundational, inspiring numerous variants such as UNet++ [15] and the DeepLab series [15, 16], which introduced atrous convolutions and spatial pyramid pooling to handle scale variations.
-Binary change detection (BCD) focuses on identifying areas of significant difference between two co-registered remote sensing images acquired at different times [17].
-Traditional BCD methods include image differencing, principal component analysis (PCA)-based techniques [18], and change vector analysis (CVA) [19].
-Deep learning has brought substantial advancements, with Siamese networks being a popular choice for learning discriminative features from image pairs [20, 21].
-These networks typically process the bi-temporal images through shared-weight encoders and then compare the resulting feature maps to identify changes.
-U-Net-like architectures have also been adapted for BCD, often by modifying the input layers to accept concatenated bi-temporal images or difference images, and training the network to output a binary change map [22].
-Semantic change detection (SCD), also known as ”from-to” change detection, provides a more detailed analysis by not only identifying where changes have occurred but also classifying the type of land cover transition (e.g., forest to urban) [23].
-This task is inherently more complex than BCD.
-A common traditional approach is post-classification comparison (PCC), where SS is performed independently on multi-temporal images, and the resulting maps are compared [24].
-However, PCC suffers from error accumulation from the individual classification steps.
-More integrated methods aim to perform simultaneous segmentation and change identification.
-Deep learning models for SCD often employ dual-branch or multi-task learning frameworks, where separate decoders might be used for segmentation at each time step and for the final change map [25].
-Critically analyzing these three tasks reveals profound interrelations and shared challenges.
-SCD inherently builds upon the principles of SS, as identifying ”from-to” transitions requires accurate semantic understanding of the landscape at different time points.
-BCD can be viewed as a simplified abstraction of SCD, where all types of semantic changes are aggregated into a binary ”change” or ”no change” outcome.
-Thus, accurate SS is a prerequisite for high-quality SCD, and insights from BCD can potentially guide the focus of SCD algorithms.
-The complementary nature of these tasks is evident: SS provides the static land cover context, BCD highlights dynamic regions, and SCD offers a comprehensive narrative of landscape transformation.
-Information from one task can, in principle, regularize or enhance the performance of others; for instance, knowing the semantic class of a pixel can refine change boundaries, or detecting a change can prompt a more detailed semantic analysis.
-Despite these intrinsic relationships and the potential for synergistic modeling, there is currently a significant lack of models or frameworks that can effectively unify semantic segmentation, binary change detection, and semantic change detection within a single, cohesive architecture.
-II-B Data Dimensionality in Remote Sensing Dense Prediction
-Remote sensing datasets utilized for dense prediction tasks are characterized by extensive variability across temporal, spectral, and spatial dimensions.
+II-A Data Dimensionality in Remote Sensing Dense Prediction
+Remote sensing datasets utilized for dense prediction tasks are characterized by extensive variability across spatial, temporal and spectral dimensions.
 This heterogeneity stems from the diverse array of satellite and airborne sensors, each with unique acquisition parameters, mission objectives, and coverage patterns, tailored for different application scenarios and geographical regions [16].
 Such variability presents a formidable challenge for developing universally applicable and robust dense prediction models.
+Remote sensing imagery exhibits substantial spatial heterogeneity stemming from two key properties: ground sampling distance (GSD) and geospatial extent.
+The GSD is sensor-dependent, ranging from sub-meter resolutions ideal for detailed urban analysis [17] to coarser resolutions suited for regional land cover mapping [18].
+The geospatial extent, meanwhile, is determined by the specific application, leading to variations in the captured area.
+Together, these properties dictate not only the pixel dimensions of an image but also the representation of its content.
+This variability poses a critical generalization challenge for deep learning models.
+A model trained for a specific resolution and image size typically suffers a significant performance degradation when applied to data with different spatial characteristics.
+This failure arises from two coupled effects: variations in GSD alter the perceived scale and texture of ground objects, while changes in geospatial extent modify the available contextual information.
 The temporal dimension in remote sensing data exhibits significant disparities.
 Revisit frequency, a critical factor for monitoring dynamic phenomena, ranges from multiple observations per day with sensors like MODIS [19], to several days (e.g., Sentinel-1 and Sentinel-2, with 5-12 day repeat cycles depending on latitude and constellation status [20, 21]), to 16 days for Landsat missions [22].
-Consequently, the temporal length of image sequences available for analysis can vary from bi-temporal pairs, commonly used in BCD and SCD [[20], [23]], to dense time series comprising hundreds of observations, which are invaluable for applications like agricultural monitoring [31].
-Models trained on data with a specific temporal frequency or sequence length may not generalize well to datasets with different temporal characteristics without substantial retraining or adaptation.
+Consequently, the temporal length of image sequences available for analysis can vary from bi-temporal pairs, commonly used in building change detection [23, 24], to dense time series comprising hundreds of observations, which are invaluable for applications like agricultural monitoring [25], vegetation forecasting [26] and moving object detection [27].
+Models trained on data with a specific temporal sequence length may not apply to datasets with different temporal characteristics without substantial retraining and adaptation.
 Spectral dimensionality is another source of major variation.
@@
 Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths.
-This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures.
+This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures [30].
 However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [31].
-Spatial characteristics, primarily ground sampling distance (GSD) or resolution, also vary widely.
-Data can range from very high resolution (VHR), with GSDs of less than 1 meter, to high resolution (HR, 1-10 meters, e.g., Sentinel-2, PlanetScope), to lower resolution.
-VHR imagery is suitable for detailed urban mapping or object detection [35], while medium to low-resolution data is often used for regional or global land cover mapping and monitoring [36].
-The spatial extent of individual scenes also differs, impacting the scale of analysis.
-Models trained on imagery of a particular GSD often struggle when applied to data with significantly different spatial resolutions, as object appearance and contextual cues change drastically.
-The inherent heterogeneity in temporal frequency, spectral composition, and spatial resolution across remote sensing datasets thus poses a substantial hurdle for developing universally applicable dense prediction models.
-Consequently, there is a pressing research gap concerning the development of adaptive model architectures or unified data processing strategies that can effectively ingest and interpret such diverse time-spectrum-space data formats at the input level for comprehensive and robust dense prediction.
+A series of models have been proposed to use hypernetworks to unify the inputs with different spectral bands, but ignored the unification of the output spectral bands [32, 33].
+The inherent heterogeneity in image size, temporal frequency, spectral composition across remote sensing datasets poses a substantial hurdle for developing universally applicable dense prediction models.
+Consequently, there is a pressing research gap concerning the development of adaptive model architectures or unified data processing strategies that can effectively ingest and interpret such diverse time-spectrum-space data formats at the input and output level for comprehensive and robust dense prediction.
+II-B Task Unification in Remote Sensing Dense Prediction
+Dense prediction in remote sensing encompasses a range of pixel-level interpretation tasks that are crucial for remote sensing applications such as environmental monitoring, urban planning, and disaster assessment [9].
+Among these, three tasks are fundamental: semantic segmentation, binary change detection, and semantic change detection.
+Semantic segmentation aims to assign a specific class label to every pixel in the satellite images, producing a detailed land cover map [34].
+Binary change detection, conversely, utilizes multi-temporal images to identify pixels where any form of change has occurred, outputting a binary map of change, no change [10].
+Bridging these two is semantic change detection, which not only detects changes across multi-temporal images but also identifies the ”from-to” nature of the change, such as a ’forest’ pixel becoming a ’building’ pixel [35].
+Given the significant conceptual overlap and inherent correlations among these tasks, a growing trend in the community is task unification through multi-task learning frameworks.
+Developing unified models to handle multiple dense prediction tasks simultaneously can lead to improved performance and efficiency [35, 13], which leverage shared representations to allow complementary information from one task to benefit others.
+A model trained jointly for multiple tasks can learn more robust feature extractors than a model trained on either task alone [36].
+For instance, SFCCD features task-specific branches for building semantic segmentation and change detection [35].
+It leverages paired data from both tasks for training, resulting in superior change detection performance compared to training the change detection branch solely with change detection data.
+Similarly, FCCDN incorporates multi-task branches for semantic segmentation and change detection, and it utilizes the weakly supervised results from the semantic segmentation branch to enhance change detection effectiveness [13].
+The core benefit of task unification is the potential to create more powerful and generalizable models while reducing the need to develop and deploy multiple specialized networks.
+Despite their promise, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness.
+The first issue is a pervasive dependence on paired data.
+Current models typically require that the training datasets are fully annotated for all constituent tasks, which means the exact same set of images must possess corresponding pixel-level labels for semantic segmentation, binary change, and semantic change.
+This stringent requirement drastically limits the pool of usable training data, as such comprehensively annotated, multi-task datasets are exceedingly rare and expensive to create.
+The second issue is the common architectural choice of using task-specific decoders.
+Many unified models employ a shared encoder to extract features but diverge into separate, specialized decoder heads for each task.
+This design can create information bottlenecks and prevent the model from fully exploiting the synergies and complementarities among the tasks at the deepest levels of feature decoding and synthesis.
+Consequently, there is an absence of a task unified dense prediction model that can effectively unify multiple core tasks through a joint modeling paradigm that operates on non-paired data.
+II-C Flexible Class Sets in Remote Sensing Dense Prediction
+The objective of most remote sensing dense prediction tasks is to categorize ground objects according to a single predefined semantic class set.
+However, the definition of this class set is not universal, which is highly dependent on the specific application and the characteristics of the geographic scene.
+This variability is evident across different tasks and scenarios.
+For example, a simple building extraction task may only require a binary class set of building, background, while binary change detection operates on change, no change.
+In contrast, a standard Land Use/Land Cover (LULC) classification task might involve a more complex set, such as water, forest, cropland, urban, barren.
+This variability is further compounded by scene-driven factors.
+The LULC class schema for a dense urban environment might need to include specific categories like commercial building, residential building, road, playground, vehicle, which would be irrelevant in a forest monitoring application.
+The latter might instead require a fine-grained class set like coniferous forest, bamboo forest, tea plantation, shrubland.
+While existing deep learning models have achieved impressive performance, they are almost universally designed for a specific and fixed class schema [15, 12, 14].
+The primary limitation of these models is their inherent inflexibility with respect to the semantic class set.
+They are designed, trained, and optimized for a fixed vocabulary of predefined classes.
+The architecture of the model, particularly the final classification layer, is hard-wired to the number and identity of these classes.
+As a result, even a minor modification to the class set would render the pretrained model unusable for the new task, such as adding a ’wetland’ category to a LULC model or splitting the ’building’ class into ’residential’ and ’commercial’.
+Adapting the model requires retraining or fine-tuning on a new dataset that includes the modified class set.
+Therefore, there lacks a unified remote sensing dense prediction model capable of flexibly adapting to diverse class sets across different tasks and scenes without necessitating costly and time-consuming retraining.
 III Methodology
 III-A Problem Formulation
-Dense prediction tasks in remote sensing generally refer to performing structured semantic inference on input images at the pixel level.
-While these tasks vary in nature, their core objective remains consistent: to extract rich semantic information from remote sensing data that includes temporal, spectral, and spatial dimensions, and to output structured prediction results.
-This paper addresses three typical tasks—semantic segmentation, semantic change detection, and binary change detection—and formulates them as a unified tensor mapping problem.
-Let the input remote sensing image be represented as a four-dimensional tensor $X\in\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}$, where $T_{1}$ denotes the length of the time sequence, $C_{1}$ represents the number of spectral channels, and $H_{1}$ and $W_{1}$ denote the height and width of the image, respectively.
-The model output is represented as a tensor $Y\in\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}}$, where $T_{2}$ represents the output time dimension, $C_{2}$ denotes the number of semantic categories, and $H_{2}$ and $W_{2}$ represent the spatial dimensions of the output image.
-In most cases, the input and output are spatially aligned, i.e., $H_{1}=H_{2}$ and $W_{1}=W_{2}$.
-Based on this representation, the three tasks considered in this paper can be uniformly described as follows:
-  • Semantic Segmentation: The model performs pixel-level multi-class classification on a remote sensing image at a given time point.
-This corresponds to the setup where $T_{1}=1$, $T_{2}=1$, and $C_{2}\geq 2$.
-  • Semantic Change Detection: The model predicts semantic labels for each time point to analyze semantic changes between any two time points. This corresponds to $T_{2}=T_{1}$, with $C_{2}\geq 2$.
-  • Binary Change Detection: The model detects changes between consecutive time points without distinguishing specific categories of change. This corresponds to $T_{2}=T_{1}-1$ and $C_{2}=2$.
-It is important to note that the input data may differ significantly in terms of time length $T_{1}$, spectral channels $C_{1}$, and spatial scale $(H_{1},W_{1})$, depending on the task type or dataset.
-To address this variability, we propose a unified modeling framework that aims to learn a mapping function $f_{\theta}$, such that
-$f_{\theta}:\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}\rightarrow\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}},$ (1)
-which can adapt to any input configuration and output dense prediction results that meet the requirements of the target task.
-This abstract modeling provides the mathematical foundation for the modular design of the subsequent methods.
-III-B Overview of the Temporal-Spectral-Spatial Unified Network (TSSUN)
-The proposed Temporal-Spectral-Spatial Unified Network (TSSUN) is designed as a unified modeling framework capable of addressing diverse dense prediction tasks across varying data modalities and spatial resolutions, as illustrated in Figure 2.
-TSSUN employs the Temporal-Spectral-Spatial Unified Strategy (TSSUS), which systematically integrates input alignment, feature extraction, and output decoding into a cohesive processing pipeline.
-In the input stage, TSSUN introduces the Spectral-Spatial Unified Module (SSUM), which encodes heterogeneous spectral and spatial data into a unified representation.
-By leveraging spectral and spatial priors through a guided alignment mechanism, SSUM ensures data consistency across spectral and spatial dimensions, establishing a coherent input foundation for subsequent feature modeling.
-During the feature extraction stage, TSSUN employs the Local-Global Window Attention (LGWA) mechanism as the primary feature modeling unit, effectively capturing multi-scale features in TSS inputs.
-LGWA integrates three distinct window-based attention blocks, each tailored to capture local and global features at different scales.
-This structure balances computational efficiency and representational power, facilitating the extraction of spatial structures across multiple scales.
-At the encoder-decoder junction, TSSUN incorporates the Temporal Unification Module (TUM), which focuses on temporal feature fusion.
-TUM mitigates temporal redundancy and adjusts the temporal length of the output based on task-specific requirements, ensuring temporal coherence in the generated predictions.
-In the output stage, TSSUN reintroduces SSUM in a mirrored configuration to further reinforce spectral-spatial consistency during the decoding process.
-By progressively decoding spectral and spatial dimensions, SSUM enables unified modeling for both semantic segmentation with arbitrary class numbers and multi-resolution spatial reconstruction, thereby enhancing TSSUN’s capability to handle multi-task, multi-modal data effectively.
-Overall, TSSUN leverages TSSUS to comprehensively integrate the input, feature extraction, and output stages, establishing a unified framework for heterogeneous data representation and dense prediction.
-The LGWA module further enhances the model’s capacity to effectively capture both local and global contextual information, ensuring a balanced representation of multi-scale and multi-modal features.
-The following sections provide a detailed analysis of the two core components, TSSUS and LGWA, including their structural design and functional mechanisms.
-Figure 2: Overview of the TSSUN architecture, comprising the Spectral-Spatial Unified Module (SSUM) for input alignment, Local-Global Window Attention (LGWA) blocks for feature extraction, and the Temporal Unified Module (TUM) for temporal fusion and output adaptation.
-The top right panel details the Dimension Unified Module (DUM), while the bottom right panel illustrates the LGWA block structure.
-III-C Temporal-Spectral-Spatial Unified Strategy (TSSUS)
-The Temporal-Spectral-Spatial Unified Strategy (TSSUS) is designed to decouple and unify the representation of inputs and outputs across temporal, spectral, and spatial dimensions.
-In the input stage, TSSUS leverages the inherent characteristics of spectral and spatial dimensions to achieve unified encoding of remote sensing imagery.
-Specifically, in the spectral dimension, TSSUS captures the continuity among spectral bands and incorporates spectral metadata to ensure spectral alignment.
-In the spatial dimension, TSSUS extracts spatial scale representations by modeling the similarity among neighboring regions and incorporating spatial metadata.
-By retaining the temporal dimension at the input stage, TSSUS enables independent feature extraction for remote sensing images across different time instances.
-At the encoder-decoder interface, TSSUS performs feature-level fusion along the temporal axis, effectively suppressing irrelevant information while adaptively adjusting the output temporal length to accommodate dense prediction tasks.
-In the output stage, TSSUS further leverages the structured characteristics of dense prediction tasks in remote sensing across spectral and spatial dimensions.
-In the spectral dimension, it models inter-class relationships to facilitate predictions with arbitrary class numbers.
-In the spatial dimension, it enhances the model’s generalization capability for multi-resolution data outputs by leveraging the spatial structure consistency among neighboring regions.
-To facilitate the mapping from the original data space to the unified feature space across different variables (temporal, spectral, and spatial), TSSUS introduces the Dimension Unified Module (DUM), which serves as the foundational implementation for both the SSUM and the TUM.
-DUM employs metadata from TSS variable subsets to generate adaptive weights and biases for linear layers, thereby enabling adaptive feature mapping.
-As illustrated in the top-right subfigure of Figure 2, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, where $D_{1}$ varies across different tasks and scenarios, DUM maps the input tensor to a unified feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process.
+Dense prediction in remote sensing encompasses a range of tasks that, despite their distinct objectives, share a common foundation: inferring structured, pixel-wise semantic information from multi-dimensional earth observation data.
+We formalize these tasks within a unified mathematical framework, conceptualizing them as a highly flexible tensor-to-tensor mapping problem.
+This abstraction is essential for accommodating the inherent heterogeneity of remote sensing data in the spatial, temporal, and spectral dimensions.
+Let an input remote sensing data instance be represented by a primary data tensor $X\in\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}$, while the corresponding model output is a prediction tensor $Y\in\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}}$, which is the same as the formulation in Section I.
+$Y$ are typically aligned with the $X$ by formulation ($H_{2}=H_{1},W_{2}=W_{1}$) and specific dense prediction task, where the spatial, temporal, and spectral dimensions of $X$ and $Y$ can vary significantly, while $T_{1}$ and $T_{2}$ is strongly correlated with the dense prediction tasks, and $C_{2}$ is strongly correlated with the predicted set of semantic categories.
+To address the profound variability in data characteristics, task and category requirements, we augment this core tensor representation with explicit metadata.
+The model’s behavior is conditioned not only on the data tensor $X$ but also on two metadata sets: an input set $M_{in}$ describing the source data, and an output set $M_{out}$ specifying the target prediction structure.
+These sets are decomposed as follows:
+  • Input Dimension Metadata: $M_{in}=\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in}\}$.
+This set provides essential context about the input data $X$.
+$M_{spa}^{in}$ encodes spatial information of input data, including pixel locations and spatial resolution.
+$M_{tem}^{in}$ contains the acquisition timestamps for each of the $T_{1}$ temporal slices.
+$M_{spe}^{in}$ specifies the wavelength for each of the $C_{1}$ spectral channels.
+  • Output Dimension Metadata: $M_{out}=\{M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$.
+This set instructs the model on the desired output format and task.
+$M_{spa}^{out}$ defines the pixel locations and spatial resolution of the prediction $Y$.
+$M_{tem}^{out}$ determines the temporal nature of the task.
+It includes the target temporal length $T_{2}$ and a trainable task embedding that specifies which dense prediction task to perform.
+$M_{spe}^{out}$ defines the prediction’s semantic space.
+It consists of a dynamically selectable subset of trainable embeddings corresponding to the $C_{2}$ target classes.
+Within this framework, our goal is to learn a single, unified mapping function $f_{\theta}$ parameterized by $\theta$, which can adapt to arbitrary input and output configurations.
+The general mapping is defined as:
+$Y=f_{\theta}(X,M_{in},M_{out}).$ (1)
+The three core tasks of semantic segmentation, binary change detection, and semantic change detection and various semantic category sets are thus handled as specific instances of this general function.
+This metadata-driven formulation provides the mathematical basis for a model that is not constrained to fixed data structures, predefined tasks or predefined category sets, enabling a truly unified approach to dense prediction in remote sensing.
+III-B Overview of the Spatial-Temporal-Spectral Unified Network
+Figure 2: Overview of the STSUN architecture, comprising the Spectral-Spatial Unified Module for input and output alignment, the Basic blocks for feature extraction, and the Temporal Unified Module for temporal fusion and output adaptation.
+The Spatial-Temporal-Spectral Unified Network unifies data representation by leveraging metadata from input and output data across spatial, temporal, and spectral dimensions.
+It employs trainable task embeddings and class embeddings to specify the particular dense prediction task and the semantic classes set.
+This design enables STSUN to adapt to various input and output shape configurations and unifies semantic segmentation, binary change detection, and semantic change detection tasks, supporting predictions across diverse class sets.
+STSUN primarily consists of five stages, as illustrated in Figure 2.
+First, at the input stage, the spatial-spectral unified module (SSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding.
+The temporal dimension, being intrinsically linked to the specific dense prediction task, necessitates a distinct processing approach from the spatial and spectral dimensions.
+If the temporal dimension were to be unified at this stage, it would be equivalent to a pixel-level fusion of multi-temporal remote sensing images.
+This would introduce significant interference from irrelevant information, thereby degrading the model’s performance on dense prediction tasks [13, 12].
+Therefore, at this stage, the independence of the temporal dimension is preserved while the spatial and spectral dimensions are unified.
+In the spatial dimension, the spatial resolutions of the input images and the spatial position of each pixel are utilized as metadata for the input spatial dimension.
+This enables a uniform representation of data based on the similarity of spatial proximity, accommodating varied resolutions.
+Concurrently, in the channel dimension, the spectral wavelength of each spectral band is incorporated as metadata for the input channel dimension, facilitating a unified data representation grounded in spectral continuity.
+Second, at the encoder stage, the encoder local global blocks are used to extract features from input data.
+As the temporal independence of remote sensing images is explicitly preserved, a shared-weight encoder then processes each temporal remote sensing image independently, extracting local and global features across distinct temporal instances within encoder local global blocks.
+Third, at the encoder–decoder junction, the temporal unified module (TUM) utilizes the input temporal metadata $M_{tem}^{in}$ to fuse features from different temporal instances and adjusts these features using the output temporal metadata $M_{tem}^{out}$ to adapt them for the specific dense prediction task.
+The temporal position within the remote sensing image time series is used as metadata for the input time dimension, which enables feature-level fusion of remote sensing image features across various time points, effectively mitigating interference from irrelevant temporal information.
+Simultaneously, to guide the dense prediction task, multiple predefined trainable task embeddings are introduced.
+A specific task embedding is selected and incorporated as metadata with output temporal length for the output time dimension, thereby instructing the model on the required dense prediction task.
+It is crucial to note that operations performed in this part maintain temporal independence and do not merge with the channel dimension [13, 12].
+This approach allows different dense prediction tasks to be unified within an independent temporal dimension, preventing interference from the remote sensing image features.
+Instead, it guides the subsequent decoder to optimize these features for the specified task, which is conducive to the model’s ability to jointly model multiple dense prediction tasks.
+Fourth, the decoder subsequently processes the fused multi-temporal remote sensing image features.
+Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder optimizes the features for the specific task by removing irrelevant information, progressively aligning them with the target output.
+Fifth, in the output stage, the SSUM utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories.
+In the spatial dimension, the image size of the remote sensing images and the pixel position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size.
+In the channel dimension, multiple predefined trainable category embeddings are established.
+From these, a subset of category embeddings is dynamically selected and integrated as metadata for the output channel dimension, explicitly indicating the set of categories the model is instructed to predict, thus enabling flexible class-agnostic prediction.
+Through the arrangement of spatial, temporal, and spectral dimensions across these five stages, and the design of metadata for each dimension, STSUN accommodates diverse input/output dimensional configurations and various subsets of prediction classes.
+Moreover, these designs enables the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection.
+By uncovering the correlations and complementarities among these tasks, STSUN leverages their combined data to enhance its performance across multiple dense prediction challenges.
+III-C Dimension Unified Module
+Figure 3: Architecture of the proposed Dimension Unified Module.
+To facilitate the mapping from the original data space to the unified feature space across different dimensions (spatial, temporal and spectral), and transform the unified features into output of appropriate shapes according to the specific task and category requirements, STSUN introduces the Dimension Unified Module (DUM), which serves as the fundamental implementation of both SSUM and TUM.
+DUM employs metadata from STS variable subsets and trainable embeddings to generate adaptive weights and biases for linear layers, thereby enabling adaptive and metadata-aware feature transformation.
+As illustrated in the Figure 3, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, as well as trainable embeddings $E\in\mathbb{R}^{D_{1}}$ when performing on output temporal and spectral dimension, where $D_{1}$ can refer to spatial, temporal or spectral dimension and varies across different tasks and scenarios, and $M$ is an metadata of $\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in},M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$.
+DUM maps the varying input tensor to a flexible feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process.
 The detailed procedure is as follows:
-  • Metadata Embedding: Metadata $M$ is subjected to positional encoding and tokenization.
+  • Metadata Embedding: Metadata $M$ is subjected to tokenization through a linear layer and positional encoding.
 A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
-  • Cross-Variable Relationship Modeling: The token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
+  • Cross-Variable Relationship Modeling: The metadata token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
   • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$.
 The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels.
-The input $X$ is first reshaped from $(D_{1},H,W)$ to $(L,D_{1})$, where $L=H\times W$, and then processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$.
-The DUM structure not only adaptively maps any TSS variable subset to a unified feature in terms of shape, but also effectively preserves the relationships between remote sensing data in original space, thereby enhancing the model’s generalization capability across heterogeneous data.
-III-D Local-Global Window Attention (LGWA)
-Transformer networks have become a prevalent structure in remote sensing tasks due to their exceptional capacity for modeling long-range dependencies.
-However, the computational complexity of traditional transformers increases quadratically with sequence length, posing significant challenges for the efficient processing of large-scale remote sensing data.
-To address this issue, this study proposes the Local-Global Window Attention (LGWA) module, which balances computational efficiency and modeling capability by capturing fine-grained features within local windows while simultaneously incorporating global contextual information.
+The input $X$ of shape $(L,D_{1})$ can be processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$.
+Specifically, when this module is applied to the output temporal and spectral dimensions, an additional learnable embedding $E$ is fused with $X$ prior to linear layer processing, which indicate the specific dense prediction task required for the features (corresponding to the output temporal dimension) or the specific set of semantic categories to be predicted (corresponding to the output spectral dimension).
+By adaptively mapping STS variable subsets to a unified feature representation, DUM not only standardizes the feature shape but also preserves important relationships present in the original remote sensing data.
+Additionally, DUM can transform this unified feature representation into an output of a specific shape as required by the task, based on the dimensionality metadata of the output data.
+This enables STSUN to effectively support inputs and outputs with various dimensional settings, allowing it to generalize to diverse dense prediction tasks and class sets.
+III-D Local-Global Window Attention
+Global attention mechanisms have become widely adopted in remote sensing tasks due to their strong capability to model long-range dependencies.
+However, their quadratic computational complexity with respect to sequence length poses a major challenge when dealing with large-scale remote sensing data.
+Moreover, global attention primarily captures holistic contextual features, often neglecting local spatial patterns, which are equally important for dense prediction tasks.
+To address this issue, this study proposes the Local-Global Window Attention module (LGWA), which improves both computational efficiency and modeling capability by jointly capturing fine-grained local features and global contextual information.
+Figure 4: Architecture of the proposed LGWA-based Basic block, employing multiple local window to extract local feature and single global window to extract global feature.
 LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 4.
-These window shapes not only determine the feature sensitivity within each window but also directly impact computational efficiency.
-Therefore, selecting appropriate window configurations in different network blocks allows for an optimal trade-off between modeling effectiveness and computational resources.
+These configurations determine the receptive field and sensitivity of the attention operation within each window and directly influence computational cost.
+Selecting appropriate window configurations in different layers allows for an optimal trade-off between feature richness and resource efficiency.
 For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@
 The attention calculation within each window focuses on extracting fine-grained features, while different window configurations provide sensitivity to varying scales.
-Figure 3: Architecture of the proposed Local-Global Window Attention (LGWA)-based Transformer block, employing multiple window shapes to enhance feature extraction.
-Next, attention scores are computed using the scaled dot-product attention mechanism:
+The attention scores are computed using the scaled dot-product attention mechanism:
 $\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (3)
@@
 where $H_{i}=\text{Attn}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$, and $W^{O}\in\mathbb{R}^{hd_{v}\times d_{M}}$ projects the concatenated output back to the original dimension.
-By adopting this modular design, the LGWA module effectively captures fine-grained features within local windows while integrating global contextual information.
-The flexible configuration of window shapes further optimizes computational efficiency, enabling the module to handle multi-scale remote sensing scenarios effectively.
+Through this modular design, LGWA effectively captures both local structural details and global semantic context.
+The flexible design of window configurations allows the module to balance modeling power and computational efficiency, making it well-suited for multi-scale, high-resolution remote sensing applications.
 IV Experiments and Results
-To validate the adaptability of TSSUN to arbitrary Temporal-Spectral-Spatial inputs and its capability to concurrently perform semantic segmentation, semantic change detection, and binary change detection tasks with support for a variable number of output classes, we conducted experiments on a total of six datasets across building and Land Cover/Land Classification (LULC) scenarios.
-For each scenario, a unified TSSUN model was trained using the combined training sets from all datasets within that scenario, and its performance was evaluated on their respective test sets.
+To validate the adaptability of STSUN to arbitrary Spatial-Temporal-Spectral inputs and outputs, and its capability to concurrently perform semantic segmentation, binary change detection, and semantic change detection tasks with support for variable class subsets, we conducted experiments on a total of six datasets across building and Land Use/Land Cover (LULC) scenarios.
+For each scenario, a unified STSUN model was trained using the combined training sets from all datasets within that scenario, and its performance was evaluated on their respective test sets.
 In the building scenario, we selected the WHU, WHU-CD, LEVIR-CD, and TSCD datasets.
-While these datasets share the same number of channels in their input and label data, they exhibit variations in temporal and spatial dimensions, leading to differences in dense prediction task types.
-Specifically, the WHU dataset corresponds to a single-temporal building semantic segmentation task.
-The WHU-CD dataset is associated with a bi-temporal building semantic change detection task.
-The LEVIR-CD dataset is used for a bi-temporal building binary change detection task.
-The TSCD dataset pertains to a multi-temporal building binary change detection task, as summarized in Table I.
+While these datasets share the same number of channels in their input and label data, they exhibit variations in temporal and spatial dimensions, leading to differences in dense prediction task types, thus verifying that STSUN can adapt to various spatiotemporal input and output settings, and can unify semantic segmentation, binary change detection, and semantic change detection tasks.
+Specifically, the WHU dataset corresponds to a single-temporal building semantic segmentation task, the WHU-CD dataset is associated with a bi-temporal building semantic change detection task, the LEVIR-CD dataset is used for a bi-temporal building binary change detection task, and the TSCD dataset pertains to a multi-temporal building binary change detection task, as summarized in Table I.
 In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets.
-These datasets differ in their temporal, spectral dimensions for both input and label data.
-The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task, while the Dynamic EarthNet dataset is used for multi-temporal semantic change detection, semantic segmentation and binary change detection tasks, as detailed in Table I.
-The experiments across multiple datasets in the building scenario were primarily designed to verify TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spatial dimensions within this context.
-Furthermore, these experiments aimed to confirm its ability to simultaneously handle semantic segmentation, semantic change detection, and binary change detection tasks.
-The datasets in the LULC scenario were mainly used to validate TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spectral dimensions.
-Additionally, these experiments were intended to demonstrate its capacity to concurrently manage semantic segmentation, binary change detection and semantic change detection tasks while supporting a flexible number of output classes.
+These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatio-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets.
+The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal semantic segmentation, binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
 TABLE I: Brief Introduction of The Experimental Datasets.
-Name Scene Task Input T Output T Input C Output C Image Size Resolution Images
-WHU Building SS 1 1 3 1 512×512 0.3 8189
-WHU-CD Building SCD 2 2 3 1 32207×15354 0.075 1
-LEVIR-CD Building BCD 2 1 3 1 1024×1024 0.5 445
-TSCD Building BCD 4 3 3 1 256×256 0.5 2700
-LoveDA Urban LULC SS 1 1 3 7 1024×1024 0.3 5987
-DynamicEarthnet LULC SCD&SS&BCD 24 24 4 6 1024×1024 3 54750
+Name Scenario Task $T_{1}$ $T_{2}$ $C_{1}$ $C_{2}$ Image Size Resolution Images
+WHU [37] Building SS 1 1 3 1 512×512 0.3 8189
+WHU-CD [37] Building SCD 2 2 3 1 1024×1024 0.075 480
+LEVIR-CD [38] Building BCD 2 1 3 1 1024×1024 0.5 445
+TSCD [39] Building BCD 4 3 3 1 256×256 0.5 2700
+LoveDA Urban [40] LULC SS 1 1 3 7 1024×1024 0.3 5987
+Dynamic Earthnet [41] LULC SCD &SS &BCD 24 24 4 6 1024×1024 3 54750
 IV-A Datasets
-We offer a brief description of the experimental building and LULC scene datasets in Table I.
-1) Building scene datasets: The WHU Building dataset [37] is divided into two main components: one containing satellite imagery and another composed of aerial photos.
+We offer a brief description of the experimental building and LULC scenario datasets in Table I.
+1) Building scenario datasets: The WHU Building dataset [37] is divided into two main components: one containing satellite imagery and another composed of aerial photos.
 In our study, we utilize the aerial photo subset, which consists of 8,189 images.
@@
 Our experiments were conducted using the original partitioning scheme and image dimensions (512×512) as specified by the WHU dataset.
-The WHU-CD dataset [37] includes bitemporal very high-resolution (VHR) aerial images taken in 2012 and 2016, which clearly highlight major changes in building structures.
+The WHU-CD dataset [37] includes bitemporal very high-resolution (VHR) aerial images taken in 2012 and 2016, which clearly highlights major changes in building structures.
 The dataset is partitioned into non-overlapping patches of 1024×1024 pixels.
@@
 The final TSCD dataset was created through uniform cropping and partitioning of these original images and derived labels.
-2) LULC scene datasets: The LoveDA dataset [39] consists of 5,987 high-resolution optical remote sensing images (with a ground sampling distance of 0.3 m) each sized at 1024×1024 pixels.
+2) LULC scenario datasets: The LoveDA dataset [40] consists of 5,987 high-resolution optical remote sensing images (with a ground sampling distance of 0.3 m) each sized at 1024×1024 pixels.
 It covers seven land cover classes: building, road, water, barren, forest, agriculture, and background.
@@
 IV-B Baseline
-To evaluate the effectiveness of the proposed TSSUN, we conducted comparative experiments with various benchmark methods on the building and LULC scene datasets.
-The benchmark methods tested on the same dataset are based on the same splitting of the dataset and use the same data.
-On the four building scene datasets, the compared CNN-based models include FCN [13], SegNet [41], U-Net [14], PSPNet [42], HRNet [43], MA-FCN [44], Deeplabv3+ [16], ResUNet [45], MAPNet [46], D-LinkNet [47], SIINet [48], FC-EF [20], FC-Siam-Diff [20], FC-Siam-Conc [20], STANet [25], DTCDSCN [49], SNUNet [50], CDNet [51], DDCNN [52], DASNet [53], DSIFN [54], HANet [55], USSFCNet [56], and SEIFNet [57], the compared transformer-based models include Segformer [58], ChangeFormer [59] and A2Net [60], and the CNN-transformer hybrid models include BDTNet [61], TransUNet [62], CMTFNet [63], BIT [64], MTCNet [65], MSCANet [66], AMTNet-50 [67], Contrast-COUD [38] and TS-COUD [38].
+To evaluate the effectiveness of the proposed STSUN, we conducted comparative experiments with various benchmark methods on building and LULC scene datasets.
+Since STSUN is adaptable to multiple datasets, it was trained and validated across all datasets in various scenarios, whereas the benchmark methods were trained and validated on individual datasets.
+On the four building scene datasets, the compared CNN-based models include FCN [42], SegNet [43], U-Net [44], PSPNet [45], HRNet [46], MA-FCN [47], Deeplabv3+ [48], ResUNet-a [49], MAPNet [50], D-LinkNet [51], SIINet [52], FC-EF [10], FC-Siam-Diff [10], FC-Siam-Conc [10], STANet [38], DTCDSCN [53], SNUNet [54], CDNet [55], DDCNN [56], DASNet [57], DSIFN [58], HANet [59], USSFCNet [60], and SEIFNet [61], the compared transformer-based models include Segformer [62], ChangeFormer [63] and A2Net [64], and the CNN-transformer hybrid models include BDTNet [65], TransUNet [66], CMTFNet [67], BIT [15], MTCNet [68], MSCANet [69], AMTNet-50 [70], Contrast-COUD [39] and TS-COUD [39].
 On the two LULC scene datasets, the compared CNN-based models include U-Net [44], Deepabv3+ [48], DANet [71], ResUNet-a [49], DASSN [72], HCANet [73], RAANet [74], A2-FPN [75] and CAC [76], the compared transformer-based models include LANet [77], SCAttNet [78], and TSViT [79], the CNN-transformer hybrid models include SAPNet [80], SSCBNet [81], UTAE [82], A2Net [83], SCanNet [84] and TSSCD [85].
@@
 Specifically, the employed transformations were limited to horizontal/vertical flipping (probability = 0.5) and transposition (probability = 0.5).
-2) Training and Inference: The TSSUN model was implemented using PyTorch [83] and executed on a single RTX A100 GPU (80G).
+2) Training and Inference: The STSUN model was implemented using PyTorch [86] and executed on a single RTX A100 GPU (80G).
 Due to the heterogeneous image resolutions across the datasets, the batch size was set to 16 for the four building scene datasets and 4 for the two LULC scene datasets.
@@
 IV-D Ablation Study
-To ascertain the efficacy of the proposed Temporal-Spectral-Spatial Unified Strategy (TSSUS) and Local-Global Window Attention (LGWA), ablation studies were performed on the TSCD dataset.
-As comparative baselines, TSSUS was replaced with a strategy involving the direct unification of temporal, spectral, and spatial dimensions at both the input and output stages ( referred to as ”direct unification”).
+To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset.
+The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions.
+As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”).
+Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension.
+The DUM is then used to map the combined channel dimension into a unified feature.
+In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task.
 Concurrently, LGWA was substituted with standard global attention.
-This experimental design serves to highlight the distinct advantages of TSSUS in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
-TABLE II: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
-Strategy Attention F1 IoU OA
-TSSUS LGWA 66.48 49.79 98.05
-Direct Unification LGWA 63.22 46.22 94.72
-TSSUS Global Attention 65.91 49.15 97.21
-Direct Unification Global Attention 62.87 45.85 94.38
-Table 2 presents the results of these ablation studies, indicating that both TSSUS and LGWA yield superior performance compared to their respective baseline alternatives.
-Specifically, TSSUS distinguishes itself from the direct unification approach.
-It begins by unifying the spectral and spatial dimensions at the input stage.
-Then, within the encoder, it independently extracts features for each temporal instance (i.e., each remote sensing image).
-Subsequently, these temporally distinct features are integrated via feature-level fusion at the interface between the encoder and decoder.
-This methodology—processing temporal information independently before fusion—effectively minimizes interference from irrelevant pixel information within the remote sensing imagery.
-Furthermore, TSSUS dynamically adjusts the temporal length of feature fusion to align with the specific requirements of the target dense prediction task.
-Following this, it optimizes the output for each time step independently within the decoder, thereby significantly enhancing the overall model efficacy.
+This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
+TABLE II: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
+Strategy Attention F1(%) IoU(%) OA(%)
+Coupled Unification Global Attention 62.87 45.85 94.38
+Coupled Unification LGWA 63.22 46.22 94.72
+Decoupled Unification Global Attention 65.91 49.15 97.21
+Decoupled Unification LGWA 66.48 49.79 98.05
+Table II displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives.
+Furthermore, their combined use leads to even greater performance improvements.
+Specifically, the decoupled unification strategy enhances performance by preserving an independent temporal dimension.
+This allows STSUN to extract features from remote sensing images at various time steps during the encoder stage.
+Subsequently, in the feature fusion stage, feature-level fusion effectively reduces interference from irrelevant information [13].
+Moreover, when performing specific tasks, this strategy enables more effective integration with task embeddings and allows for the adjustment of temporal length to suit task requirements.
+This capability facilitates the joint modeling of multiple dense prediction tasks, leveraging data from various tasks to collectively improve their individual performance.
 In a similar vein, LGWA offers advantages over standard global self-attention.
@@
 IV-E Overall Comparison
-IV-E1 Building
-The efficacy of the proposed Temporal-Spectral-Spatial Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection.
-Our method was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
-On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed TSSUN achieves state-of-the-art performance.
-As detailed in Table III, TSSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%.
-This superior performance can be attributed to TSSUN’s Temporal-Spectral-Spatial Unified Strategy, particularly its spatial unification component, which effectively processes varying spatial resolutions, and the Local-Global Window Attention mechanism.
+IV-E1 Building Scenario
+The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection.
+STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
+Figure 5: Sample inference results on for building scene datasets.
+The input images, ground truths and predictions are shown in the first, second and third rows, respectively.
+Red areas denote false positives and blue areas denote false negatives.
+(a) WHU dataset sample.
+(b) WHU-CD dataset sample.
+(c) LEIVR-CD dataset sample.
+(d) TSCD dataset sample.
+On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed STSUN achieves state-of-the-art performance.
+As detailed in Table III, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%.
+This superior performance can be attributed to STSUN, particularly its spatial unification component, which effectively processes varying image size, and the Local-Global Window Attention mechanism.
 The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
@@
 Deeplabv3+ [48] 94.31 94.53 94.42 89.43
-ResUNet [45] 94.49 94.71 94.60 89.75
+ResUNet-a [49] 94.49 94.71 94.60 89.75
 MAP-Net [50] 93.99 94.82 94.40 89.40
@@
 CMTFNet [67] 90.12 95.21 92.59 86.21
-TSSUN 95.71 94.87 95.29 91.00
-For the bi-temporal semantic change detection task on the WHU-CD dataset, TSSUN demonstrates leading performance against other SOTA methods, as shown in Table IV.
+STSUN 95.71 94.87 95.29 91.00
+For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table IV.
 The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points.
-TSSUN excels here due to its inherent design for unified temporal and spectral modeling.
-The TSSUS component allows the network to effectively learn temporal evolutionary patterns and relationships between different land cover classes, while the LGWA mechanism enhances the discrimination of subtle yet significant semantic changes from unchanged areas, leading to more precise change maps.
+STSUN excels here due to its inherent design for unified temporal and spectral modeling.
+The STSUN component allows the network to effectively learn temporal evolutionary patterns and relationships between buildings and other objects, leading to more precise change maps.
 TABLE IV: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
@@
 TransUNet [66] 93.82 89.33 91.52 84.37
-RSM-CD 93.07 91.20 92.13 85.40
-In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, TSSUN surpasses existing methods.
-Table V shows that our method achieves the highest F1-score of 91.59% and an IoU of 84.49%.
-The strength of TSSUN on this dataset lies in its robust temporal modeling capabilities, facilitated by TSSUS, which allows for consistent feature representation across different time points.
+STSUN 93.07 91.20 92.13 85.40
+In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, STSUN surpasses compared methods.
+Table V shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%.
+The strength of STSUN on this dataset lies in its robust temporal modeling capabilities, which allows for consistent feature representation across different time points.
 Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
@@
 AMTNet-50 [70] 91.82 89.71 90.76 83.08
-TSSUN 93.17 90.07 91.59 84.49
-Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, TSSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%.
+STSUN 93.17 90.07 91.59 84.49
+Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%.
 The TSCD dataset requires robust modeling of temporal dependencies across multiple observations.
-TSSUS is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of land cover.
-This, combined with the LGWA’s capacity to aggregate contextual information effectively over extended temporal sequences, enables TSSUN to accurately identify changes in complex, evolving landscapes where other methods might falter due to the increased temporal dimensionality.
+STSUN is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of buildings, which enables STSUN to accurately identify changes in complex, evolving landscapes.
 TABLE VI: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
@@
 TS-COUD [39] 65.24 49.33 97.90
-TSSUN 66.48 49.79 98.05
-Figure 4: Sample inference results on for building scene datasets.
-The input images, ground truths and predictions are shown in the first, second and third rows, respectively.
-Red areas denote false positives and blue areas denote false negatives.
-(a) WHU dataset sample.
-(b) WHU-CD dataset sample.
-(c) LEIVR-CD dataset sample.
-(d) TSCD dataset sample.
-Figure 4 presents a qualitative comparison of our method against other SOTA approaches on representative samples from all four datasets.
-Visually, TSSUN consistently produces more accurate and complete segmentation and change maps.
+STSUN 66.48 49.79 98.05
+Figure 5 presents inference results of STSUN from all four datasets.
+Visually, STSUN consistently produces accurate and complete segmentation and change maps.
 The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations.
-For instance, in building segmentation on the WHU dataset (Figure 4(a)), TSSUN generates sharper edges and more complete building shapes.
-Similarly, for change detection tasks on WHU-CD (Figure 4(b)), LEVIR-CD (Figure 4(c)), and TSCD (Figure 4(d)), our method demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas.
-This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of TSSUS and LGWA, to learn highly discriminative features and effectively model contextual relationships across the temporal, spectral, and spatial dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
-IV-E2 LULC
-The efficacy of the proposed Temporal-Spectral-Spatial Unified Network is rigorously evaluated against SOTA methods on two challenging benchmark datasets: LoveDA for single-temporal semantic segmentation and DynamicEarthNet for multi-temporal land cover classification.
-On the LoveDA dataset, TSSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII.
-This leading performance can be attributed to TSSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information.
+For instance, in building segmentation on the WHU dataset (Figure 5(a)), STSUN generates sharp edges and complete building shapes.
+Similarly, for change detection tasks on WHU-CD (Figure 5(b)), LEVIR-CD (Figure 5(c)), and TSCD (Figure 5(d)), STSUN demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas.
+This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of STSUN and LGWA, to learn highly discriminative features and effectively model contextual relationships across the spatial, temporal and spectral dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
+IV-E2 LULC Scenario
+The efficacy of the proposed Spatial-Temporal-Spectral Unified Network is rigorously evaluated against SOTA methods on two challenging LULC datasets: LoveDA for single-temporal land cover classification and DynamicEarthNet for multi-temporal land cover classification.
+On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII.
+This leading performance can be attributed to STSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information.
 This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery.
-Furthermore, the Temporal-Spectral-Spatial Unified Strategy enables robust feature representation by adeptly unifying spatial and spectral information, which is vital for handling the heterogeneity present in such diverse urban and rural landscapes with uneven class distributions.
-For the multi-temporal DynamicEarthNet dataset, TSSUN again surpasses existing methods, yielding the top scores across all reported metrics: Semantic Change Segmentation (SCS) score of 29.9, class-agnostic Binary Change score (BC) of 38.9, and mIoU of 54.7 as shown in Table VIII.
-The success of TSSUN on this dataset underscores the effectiveness of the TSSUS in explicitly modeling the temporal dimension.
-By leveraging temporal meta-information to ensure consistency across sequences of varying lengths, TSSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
-This robust temporal modeling, distinct from its spatial feature extraction strengths highlighted in the LoveDA analysis, allows for consistent and accurate classification and change analysis across the 24 annotated time steps.
 TABLE VII: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
-Methods F1-score per category (%) AF OA mIoU
+Methods F1-score per category (%) AF(%) OA(%) mIoU(%)
 Background Building Road Water Barren Forest Agriculture
@@
 SSCBNet [81] 68.25 75.90 79.00 89.10 54.00 66.20 82.15 74.05 70.95 64.58
-TSSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
+STSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
+For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VIII.
+The success of STSUN on this dataset underscores the effectiveness of the STSUN in explicitly modeling the temporal dimension.
+By leveraging temporal metadata to ensure consistency across sequences of varying lengths, STSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
 TABLE VIII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.
@@
 TSSCD [85] 12.0 19.4 4.7 33.9
-TSSUN 29.9 38.9 21.8 54.7
+STSUN 29.9 38.9 21.8 55.7
+Figure 6 presents the LULC semantic segmentation results of our proposed STSUN on the single-temporal LoveDA dataset (Figure 6 a) and the multi-temporal Dynamic EarthNet dataset (Figure 6 b).
+The method achieves strong performance across both, effectively navigating challenges such as the complex spectral-spatial features, multi-scale objects in LoveDA, and the inherent temporal variations within Dynamic EarthNet.
+This proficiency stems from STSUN’s ability to perform unified representation and modeling of remote sensing data across spatial, temporal and spectral dimensions.
+Specifically, the STSUN enables harmonized encoding and feature fusion across these dimensions, while the LGWA mechanism efficiently captures both local details and global contextual information, crucial for accurate LULC delineation in diverse scenarios.
+Figure 6: Sample inference results on for LULC scene datasets.
+The input images, ground truths and predictions are shown in the first, second and third rows, respectively.
+(a) LoveDA dataset sample.
+(b) DynamicEarthNet dataset sample.
 V DISCUSSION
-The proliferation of diverse remote sensing data sources necessitates models capable of unified processing across varied data characteristics and task requirements.
-This research confronts the critical challenge of existing deep learning models in remote sensing dense prediction, which are typically constrained to fixed input-output configurations and struggle to generalize across the inherent heterogeneity in temporal length, spectral channels, and spatial resolution of imagery, as well as diverse task specifications such as semantic segmentation, semantic change detection, and binary change detection.
-This rigidity imposes significant burdens in terms of retraining and fine-tuning for each specific scenario, limiting practical applicability in dynamic operational environments.
-To surmount this limitation, we proposed the Temporal-Spectral-Spatial Unified Network, a novel architecture engineered for comprehensive unification.
-The cornerstone of TSSUN is the Temporal-Spectral-Spatial Unified Strategy, which systematically decouples and standardizes data representations.
-On the input side, TSSUS leverages dimension-specific meta-information to harmonize temporal sequences of varying lengths, align spectral inputs with differing channel counts, and reconcile features from disparate spatial resolutions.
-This is achieved by modeling the continuity in temporal evolution, the inherent relationships between spectral bands, and the local similarity in spatial regions.
-Concurrently, on the output side, TSSUS exploits shared structural elements among the three primary dense prediction tasks.
-It supports variable temporal output lengths by identifying common temporal patterns, accommodates an arbitrary number of output classes by capturing inter-class relationships, and ensures spatial consistency across resolutions.
-Furthermore, the integration of a Local-Global Window Attention mechanism enhances TSSUN’s capacity.
-LGWA efficiently processes local contextual information through overlapping windowed attention blocks and aggregates global context via a subsequent global attention layer, striking a balance between computational load and feature representation power essential for complex remote sensing scenes.
-The efficacy of the TSSUN framework in addressing the identified research problem is substantiated by extensive experimental validation.
-Results across multiple datasets, encompassing diverse Temporal-Spectral-Spatial configurations and tasks like building detection and land use/land cover classification, demonstrate that a single, pre-trained TSSUN model can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks, consistently achieving performance comparable to or exceeding state-of-the-art specialized models.
-Despite these advancements, the current TSSUN framework possesses certain limitations.
-Firstly, the TSSUS component relies on the availability and accuracy of explicit meta-information (e.g., acquisition dates, sensor band specifications, ground sampling distance) to guide the unification across temporal, spectral, and spatial dimensions.
-The performance may be suboptimal if such meta-information is imprecise, incomplete, or unavailable, potentially hindering the model’s adaptability in data-scarce or poorly documented scenarios.
-Secondly, while the LGWA mechanism is designed for efficiency, the computational demand of its global attention component, although optimized, might still present a scalability challenge when processing extremely large-scale images or exceptionally long and dense temporal sequences, potentially requiring substantial memory and processing resources.
-Future research will focus on mitigating these limitations and expanding the capabilities of TSSUN.
-One promising direction involves developing methods for the implicit learning or unsupervised inference of the requisite meta-information directly from the remote sensing data, thereby reducing dependency on external inputs and enhancing robustness.
-Another avenue for exploration is the refinement of the LGWA module, possibly by incorporating more sophisticated hierarchical attention structures or adaptive computational pruning techniques to further improve its efficiency and scalability for massive datasets.
-Additionally, future work could extend the unification paradigm of TSSUN to encompass a broader array of remote sensing tasks and potentially integrate other data modalities, such as SAR or LiDAR, within a unified analytical framework.
+V-A Unified Multitask Learning
+STSUN introduces a unified framework for dense prediction, capable of concurrently addressing semantic segmentation, binary change detection, and semantic change detection within a single model.
+This unification is principally achieved through the trainable task embeddings and the Dimension Unification Module.
+This approach aligns with the core tenets of multi-task learning (MTL), which posit that jointly learning related tasks can lead to improved generalization by leveraging shared representations [90].
+By training on a comprehensive dataset aggregated from these distinct tasks, our model develops more robust features than models trained in a single-task paradigm [91].
+Moreover, our framework marks a departure from conventional MTL applications in the remote sensing domain.
+Prevailing methods typically depend on the availability of paired datasets, where each geographical location has labels for all tasks, and commonly employ multiple task-specific decoder heads to generate predictions.
+In stark contrast, the key strengths of our model are its capacity to be trained on non-paired data from disparate tasks and its reliance on a single, shared decoder head for all predictions.
+The implications of this design are twofold.
+First, the ability to utilize non-paired data dramatically expands the pool of usable training imagery, addressing the data scarcity issue in the field.
+Second, the single shared decoder enforces the learning of a more cohesive and generalized feature representation, promoting more effective knowledge transfer between tasks.
+This results in a framework that is not only more data-efficient but also exhibits enhanced adaptability and robustness across a diverse range of dense prediction challenges.
+V-B Limitation and Expectation
+The increasing heterogeneity of remote sensing data across spatial, temporal, and spectral dimensions presents a fundamental challenge to conventional deep learning models, which are typically engineered for fixed data configurations and singular tasks.
+This rigidity impedes the integrated analysis of multi-source data, such as fusing Landsat and Sentinel imagery for longitudinal land cover analysis or adapting a single model for both routine monitoring and rapid disaster assessment.
+Our work addresses this critical gap by introducing a unified framework, STSUN, which demonstrates that a single model can adeptly navigate these STS variations.
+The scientific significance of this approach lies in its ability to holistically leverage the vast and disparate archives of remote sensing data, fostering the development of more robust, flexible, and data-efficient solutions for dense prediction.
+Our experimental results robustly demonstrate that a single STSUN can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks.
+Across multiple datasets, STSUN consistently achieved performance comparable to or exceeding that of state-of-the-art specialized models, highlighting its significant potential for streamlining complex remote sensing data analysis workflows and reducing the need for task-specific model development.
+The superior performance of STSUN, as evidenced by our extensive experiments, stems from its departure from traditional multi-task learning paradigms.
+Prevailing MTL methods in remote sensing often require co-registered, fully labeled datasets for all tasks and rely on multiple task-specific decoders.
+In contrast, STSUN’s architecture, centered on a shared decoder modulated by task and category embeddings, facilitates joint modeling on a diverse corpus of non-paired data aggregated from semantic segmentation, binary change detection, and semantic change detection.
+Training on this wider spectrum of data forces the model to learn a more generalized and powerful feature representation by extracting shared knowledge across tasks and data sources, leading to enhanced performance without the need for task-specific architectural components.
+Despite its promising results, the current implementation of STSUN has two principal limitations that suggest clear avenues for future research.
+First, the framework’s flexibility is contingent upon explicit metadata, such as predefined task and category embeddings, which must be provided at inference time.
+A compelling future direction involves exploring methods for the model to implicitly infer the task and desired output configuration directly from the input data context, thereby increasing its autonomy.
+Second, the potential of our unified modeling approach has been demonstrated only within a supervised learning context.
+A significant opportunity exists to extend this framework to the self-supervised domain to construct vision foundation models and vision language models.
+By pre-training the vision model on massive, unlabeled remote sensing archives with diverse STS characteristics, one could develop a powerful remote sensing vision foundation model.
+Such a model, pre-trained to understand the fundamental structure of STS data, could then be fine-tuned with minimal data for a wide array of downstream applications or potentially used in zero-shot settings, drastically reducing the dependence on labeled data and advancing the capabilities of large-scale Earth observation.
+At the same time, applying this method to remote sensing vision-language models could enable the models to be trained with large amounts of remote sensing image-text pair data without needing to consider the dimensionality settings of the remote sensing images.
+This would not only enhance the performance of remote sensing vision-language models but also make them adaptable to various STS settings for remote sensing image and text pair data.
 VI CONCLUSION
-This paper addressed the critical challenge of processing heterogeneous remote sensing data for diverse dense prediction tasks, a common hurdle due to variations in temporal, spectral, and spatial characteristics.
-We introduced the Temporal-Spectral-Spatial Unified Network, a novel deep learning architecture.
-TSSUN demonstrates remarkable flexibility by handling arbitrary input configurations across these three dimensions and unifying key dense prediction tasks: semantic segmentation, semantic change detection, and binary change detection, with adaptable output class settings.
-The proposed Temporal-Spectral-Spatial Unified Strategy enables this unification by effectively decoupling and modeling data representations guided by meta-information.
-Furthermore, the designed Local-Global Window Attention mechanism significantly improves feature extraction by capturing both local and global contextual information efficiently.
-Comprehensive experimental validation on multiple datasets confirmed that a single TSSUN model consistently adapts to varied data inputs and task requirements, achieving or exceeding state-of-the-art performance.
-The TSSUN framework represents a substantial advancement towards universal remote sensing models, mitigating the need for task-specific designs and extensive retraining, thus paving the way for more scalable and cost-effective analysis of complex, multi-source geospatial data and fostering broader application of remote sensing technologies.
+This study introduced the Spatial-Temporal-Spectral Unified Network, a novel architecture designed to overcome the critical limitations of model rigidity in remote sensing dense prediction.
+By establishing a unified representation for arbitrary spatial, temporal, and spectral data configurations, STSUN demonstrates exceptional adaptability to heterogeneous datasets.
+STSUN uniquely integrates semantic segmentation, binary change detection, and semantic change detection within a single model via trainable task embeddings, obviating the need for specialized, task-specific architectures.
+Moreover, STSUN utilizes trainable semantic category embeddings to perform dense prediction for any prediction class setting without requiring model retraining.
+Comprehensive experimental validation on multiple datasets confirmed that a single STSUN model consistently adapts to varied data inputs and outputs, unifing multiple dense prediction tasks and prediction category settings, achieving or exceeding state-of-the-art performance.
+We envision STSUN to serve a baseline for universal remote sensing dense prediction models, mitigating the need for task-specific designs and extensive retraining.

### v3

@@
 Moreover, STSUN unifies disparate dense prediction tasks within a single architecture by conditioning the model on trainable task embeddings.
-Similarly, STSUN facilitates flexible prediction across any set of semantic categories by integrating trainable category embeddings as metadata.
-Extensive experiments on multiple datasets with diverse STS configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions.
+Similarly, STSUN facilitates flexible prediction across multiple set of semantic categories by integrating trainable category embeddings as metadata.
+Extensive experiments on multiple datasets with diverse Spatial-Temporal-Spectral configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions.
 The proposed approach consistently achieves state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications.
@@
 Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7].
-Remote sensing imagery exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands in practical applications [8].
+Remote sensing imagery and label exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions in dense prediction tasks, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands of input and output in practical applications [8].
 Remote sensing dense prediction refers to tasks where the goal is to produce a pixel-wise prediction, assigning a specific label to every pixel in the input satellite imagery [9].
-From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection.
-These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively.
+Dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection.
+These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, a dense prediction model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively.
 Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$.
-In semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$.
-Therefore, the temporal dimension $T_{1},T_{2}$ is strongly correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is strongly correlated with the set of predicted semantic categories.
+From the task perspective, in semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between adjacent two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$.
+Therefore, the temporal dimension $T_{1},T_{2}$ is correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is correlated with the set of predicted semantic categories.
+From the data perspective, the input’s spatial dimension is related to the geographical coverage and spatial resolution of remote sensing data; the input’s temporal dimension correlates with the temporal coverage and temporal resolution of remote sensing data; the input’s spectral dimension corresponds to the modality of remote sensing data.
+This underscores the importance of unified modeling of inputs and outputs in the spatial-temporal-spectral dimensions to achieve a unified approach for remote sensing dense prediction tasks.
 Figure 1: Illustration of Spatial-Temporal-Spectral Unified Network.
@@
 1.
-Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H1,W1),(H2,W2))$, temporal lengths $(T1,T2)$ and spectral bands $(C_{1},C_{2})$.
+Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H_{1},W_{1}),(H_{2},W_{2}))$, temporal lengths $(T_{1},T_{2})$ and spectral bands $(C_{1},C_{2})$.
 This rigidity restricts their adaptability to the diverse and heterogeneous remote sensing datasets encountered in practical applications, such as diverse satellite sensor data with varying spectral bands or multi-temporal imagery sequences with varying temporal lengths.
@@
 3.
-Flexible Category Set: STSUN facilitates flexible prediction across an arbitrary set of semantic categories, which is accomplished by using a predefined and trainable category embedding set.
+Flexible Category Set: STSUN facilitates flexible prediction across multiple sets of semantic categories, which is accomplished by using a predefined and trainable category embedding set.
 The selected category subset is integrated into the output channel dimension as metadata, which guides the model to identify the specific classes needed for a given prediction.
@@
   3. 3.
-    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for any specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
+    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for multiple specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
   4. 4.
@@
 This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures [30].
-However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [31].
+However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing spectral bandss and spectral information content [31].
 A series of models have been proposed to use hypernetworks to unify the inputs with different spectral bands, but ignored the unification of the output spectral bands [32, 33].
@@
 The core benefit of task unification is the potential to create more powerful and generalizable models while reducing the need to develop and deploy multiple specialized networks.
-Despite their promise, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness.
+Despite their superior performance, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness.
 The first issue is a pervasive dependence on paired data.
@@
 III-B Overview of the Spatial-Temporal-Spectral Unified Network
-Figure 2: Overview of the STSUN architecture, comprising the Spectral-Spatial Unified Module for input and output alignment, the Basic blocks for feature extraction, and the Temporal Unified Module for temporal fusion and output adaptation.
+Figure 2: Overview of the STSUN architecture, comprising the Input Spatial-Spectral Unified Module for spatial and spectral unification of input, the Local Global Basic blocks for global and local feature extraction, the Temporal Unified Module for temporal unification of input and output, and the Output Spatial-Spectral Unified Module for spatial and spectral unification of output.
+The highlighted portions in the shapes on either side of each unification module indicate the dimension that the module unifies.
 The Spatial-Temporal-Spectral Unified Network unifies data representation by leveraging metadata from input and output data across spatial, temporal, and spectral dimensions.
@@
 STSUN primarily consists of five stages, as illustrated in Figure 2.
-First, at the input stage, the spatial-spectral unified module (SSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding.
+First, at the input stage, the input spatial-spectral unified module (ISSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding of spatial and spectral dimensions.
 The temporal dimension, being intrinsically linked to the specific dense prediction task, necessitates a distinct processing approach from the spatial and spectral dimensions.
@@
 As the temporal independence of remote sensing images is explicitly preserved, a shared-weight encoder then processes each temporal remote sensing image independently, extracting local and global features across distinct temporal instances within Encoder Local Global Blocks.
+The overall structure of the Encoder Local Global Blocks is the same as the transformer block, except that the global attention is replaced with local-global window attention, as shown in Figure 4.
 Third, at the encoder–decoder junction, the temporal unified module (TUM) utilizes the input temporal metadata $M_{tem}^{in}$ to fuse features from different temporal instances and adjusts these features using the output temporal metadata $M_{tem}^{out}$ to adapt them for the specific dense prediction task.
@@
 Fourth, the decoder subsequently processes the fused multi-temporal remote sensing image features.
-Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder optimizes the features for the specific task by removing irrelevant information, progressively aligning them with the target output.
-Fifth, in the output stage, the SSUM utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories.
-In the spatial dimension, the image size of the remote sensing images and the pixel position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size.
+Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder local-global blocks optimize the features for the specific task by removing irrelevant information, progressively aligning them with the target output.
+The Decoder Local Global Block are the same as Encoder Local Global Block.
+Fifth, in the output stage, the output spatial-spectral unified module (OSSUM) utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories.
+In the spatial dimension, the spatial resolution of the remote sensing images and the patch position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size.
 In the channel dimension, multiple predefined trainable category embeddings are established.
@@
 Through the arrangement of spatial, temporal, and spectral dimensions across these five stages, and the design of metadata for each dimension, STSUN accommodates diverse input/output dimensional configurations and various subsets of prediction classes.
-Moreover, these designs enables the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection.
-By uncovering the correlations and complementarities among these tasks, STSUN leverages their combined data to enhance its performance across multiple dense prediction challenges.
-III-C Dimension Unified Module
-Figure 3: Architecture of the proposed Dimension Unified Module.
-To facilitate the mapping from the original data space to the unified feature space across different dimensions (spatial, temporal and spectral), and transform the unified features into output of appropriate shapes according to the specific task and category requirements, STSUN introduces the Dimension Unified Module (DUM), which serves as the fundamental implementation of both SSUM and TUM.
-DUM employs metadata from STS variable subsets and trainable embeddings to generate adaptive weights and biases for linear layers, thereby enabling adaptive and metadata-aware feature transformation.
-As illustrated in the Figure 3, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, as well as trainable embeddings $E\in\mathbb{R}^{D_{1}}$ when performing on output temporal and spectral dimension, where $D_{1}$ can refer to spatial, temporal or spectral dimension and varies across different tasks and scenarios, and $M$ is an metadata of $\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in},M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$.
-DUM maps the varying input tensor to a flexible feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process.
-The detailed procedure is as follows:
-  • Metadata Embedding: Metadata $M$ is subjected to tokenization through a linear layer and positional encoding.
-A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
-  • Cross-Variable Relationship Modeling: The metadata token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
-  • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$.
-The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels.
-The input $X$ of shape $(L,D_{1})$ can be processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$.
-Specifically, when this module is applied to the output temporal and spectral dimensions, an additional learnable embedding $E$ is fused with $X$ prior to linear layer processing, which indicate the specific dense prediction task required for the features (corresponding to the output temporal dimension) or the specific set of semantic categories to be predicted (corresponding to the output spectral dimension).
-By adaptively mapping STS variable subsets to a unified feature representation, DUM not only standardizes the feature shape but also preserves important relationships present in the original remote sensing data.
-Additionally, DUM can transform this unified feature representation into an output of a specific shape as required by the task, based on the dimensionality metadata of the output data.
-This enables STSUN to effectively support inputs and outputs with various dimensional settings, allowing it to generalize to diverse dense prediction tasks and class sets.
+Moreover, these designs enable the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection.
+By uncovering the correlations and complementarities among these tasks, STSUN leverages the combined data of multiple tasks to enhance its performance across multiple dense prediction challenges.
+III-C Spatial-Temporal-Spectral Unified Module
+Figure 3: Illustration of the Spatial-Temporal-Spectral Unified Module. (a) Input Spatial-Spectral Unified Module. (b) Temporal Unified Module. (c) Output Spatial-Spectral Unified Module.
+[…]
+The core mechanism of STSUM involves using metadata from each dimension, along with optional trainable embeddings, to generate adaptive linear layers.
+This enables the transformation of variable input data into unified features or, conversely, the conversion of unified features into variable outputs tailored to specific requirements.
+The feature mapping mechanism is consistent across the modules and comprises two main branches: a hyper-network branch and a mapping network branch.
+In the hyper-network branch, metadata from each dimension is first tokenized using a linear layer and augmented with positional encodings.
+These tokens are then processed through several Transformer blocks to capture the latent relationships among them.
+Finally, another linear layer generates the parameters for the adaptive mapping network.
+In the mapping network branch, this dynamically generated network applies a linear transformation to the input features, thereby unifying the input data or generating the adaptive output.
+Although the underlying principle is similar, each module requires distinct and meticulous design considerations regarding feature shapes and parameter generation.
+This ensures that the modules can achieve generic and robust dimensional unification by processing different metadata for different mapping requirements across various dimensions.
+III-C1 Input Spatial-Spectral Unified Module
+[…]
+[…]
+III-C2 Temporal Unified Module
+The TUM unifies the temporal dimension of the input and output data.
+It employs a similar mechanism with ISSUM to first map the variable input temporal length $T_{1}$ to a predefined, unified length $T$, and then map this unified length to a variable output length $T_{2}$ based on task requirements, as shown in Figure 3(b).
+For simplicity, let $C=C_{e}\cdot C_{a}$, $L=H\cdot W$.
+The TUM isolates the temporal dimension, obtaining a single data sample $X^{\prime}\in\mathbb{R}^{T_{1}\times L\times C}$, and reshapes it to $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{1}}$, thereby transposing the temporal dimension to the position of the channel dimension.
+This allows the TUM to unify the temporal dimension of the input data using a similar method of ISSUM.
+TUM utilizes the temporal position information of the input data as metadata, $M_{\text{tem}}^{\text{in}}\in\mathbb{R}^{T_{1}}$, to adaptively generate a linear layer composed of a bias parameter $b\in\mathbb{R}^{T}$ and a weight matrix $W\in\mathbb{R}^{T_{1}\times T}$.
+This mapping network transforms the temporally variable input $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{1}}$ into a unified representation $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T}$.
+Subsequently, according to the requirements of the specific dense prediction task, the TUM needs to map the unified temporal length $T$ to a variable length $T_{2}$.
+Unlike the previous operation of mapping variable features to unified features, this process is reversed, which necessitates differences in the metadata and the generated parameters.
+Specifically, the TUM uses the output temporal length information and a selected task embedding as the output temporal metadata, $M_{\text{tem}}^{\text{out}}\in\mathbb{R}^{T_{2}}$.
+These are mapped through linear layers and then summed to form the metadata token sequence.
+The task embedding is a predefined, trainable embedding selected from a set, such as {semantic segmentation embedding, binary change detection embedding, semantic change detection embedding}, to specify the dense prediction task being performed.
+The metadata token sequence is then processed by a Transformer block to capture latent relationships between tokens.
+Following this, it passes through a linear layer to generate only the weight parameter $W\in\mathbb{R}^{T_{2}\times T}$, which is reshaped to $W\in\mathbb{R}^{T\times T_{2}}$ to serve as the weights of the mapping network.
+Finally, this bias-free mapping network transforms $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T}$ into $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{2}}$, which is then reshaped back to $X^{\prime}\in\mathbb{R}^{T_{2}\times L\times C}$, thus converting the unified temporal length $T$ into a variable length $T_{2}$ according to specific task demands.
+The bias parameter is not generated because the mapping network would require a variable bias $b\in\mathbb{R}^{T_{2}}$, which cannot be generated from a fixed CLS token through a fixed linear layer.
+III-C3 Output Spatial-Spectral Unified Module
+The OSSUM unifies the spatial and spectral dimensions of the output data.
+It first maps the unified image size $(H,W)$ to a variable size $(H_{2},W_{2})$ using spatial metadata, and then maps the unified spectral spectral bands $C$ to a variable count $C_{2}$ using spectral metadata, as illustrated in Figure 3(c).
+The dimensional mapping mechanism of the OSSUM is similar to that of the TUM for the output temporal dimension, with the main difference lying in the metadata.
+Specifically, for the spatial dimension, the hyper-network branch of the OSSUM uses the patch position and spatial resolution as metadata, $M_{\text{spa}}^{\text{out}}\in\mathbb{R}^{P_{h}\cdot P_{w}}$.
+After passing through linear layers and Transformer blocks, it generates a weight parameter $W\in\mathbb{R}^{C_{a}\times(P_{h}\cdot P_{w})}$ for the mapping network.
+Since the output spatial size is identical to the input spatial size, we have $M_{\text{spa}}^{\text{out}}=M_{\text{spa}}^{\text{in}}$.
+In the mapping network branch, the input feature $X^{\prime}\in\mathbb{R}^{(H\cdot W)\times(C_{e}\cdot C_{a})}$ is reshaped to $X^{\prime}\in\mathbb{R}^{(H\cdot W\cdot C_{e})\times C_{a}}$.
+It is then transformed by the mapping network to yield $X^{\prime}\in\mathbb{R}^{(H\cdot W\cdot C_{e})\times(P_{h}\cdot P_{w})}$.
+An ‘unpatchify‘ operation performs up-sampling, and a final reshape operation converts it to $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{e}}$, producing a variable-sized output in the spatial dimension.
+Similarly, in the spectral dimension, the OSSUM first selects a subset from a predefined set of semantic class embeddings based on the requirements of the prediction class set.
+This subset of embeddings indicates all the classes the model needs to predict.
+For instance, while multiple remote sensing scenes might require dense prediction for various land cover types, leading to a total semantic class embedding set like {Tree, Water, Soil, Road, Building, Background}, a specific building extraction task would only require selecting the subset {Building, Background}.
+This subset directs the model to classify land cover into these two categories, enabling effective building extraction.
+Therefore, the selected subset of semantic class embeddings serves as the output spectral metadata, $M_{\text{spe}}^{\text{out}}\in\mathbb{R}^{C_{2}\times C_{e}}$.
+Through a similar mechanism above, a weight parameter $W\in\mathbb{R}^{C_{e}\times C_{2}}$ is generated for the mapping network.
+In the mapping network branch, $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{e}}$ undergoes a linear transformation by the mapping network to produce $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{2}}$.
+Finally, a reshape operation yields the output $Y\in\mathbb{R}^{H_{2}\times W_{2}\times C_{2}}$, producing a variable output in the spectral dimension.
 III-D Local-Global Window Attention
-Global attention mechanisms have become widely adopted in remote sensing tasks due to their strong capability to model long-range dependencies.
-However, their quadratic computational complexity with respect to sequence length poses a major challenge when dealing with large-scale remote sensing data.
-Moreover, global attention primarily captures holistic contextual features, often neglecting local spatial patterns, which are equally important for dense prediction tasks.
-To address this issue, this study proposes the Local-Global Window Attention module (LGWA), which improves both computational efficiency and modeling capability by jointly capturing fine-grained local features and global contextual information.
-Figure 4: Architecture of the proposed LGWA-based Basic block, employing multiple local window to extract local feature and single global window to extract global feature.
-LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 4.
-These configurations determine the receptive field and sensitivity of the attention operation within each window and directly influence computational cost.
-Selecting appropriate window configurations in different layers allows for an optimal trade-off between feature richness and resource efficiency.
+After unifying the diverse spatial-temporal-spectral input and output data, STSUN requires a powerful feature extraction capability to be effectively applied to various remote sensing scenarios and tasks.
+Therefore, this study proposes the Local-Global Window Attention module.
+This module performs self-attention operations within multiple local windows of varying shapes and a single global window, As illustrated in Figure 4.
+This design enables the joint capture of fine-grained local details and coarse-grained global features, thereby effectively extracting the local characteristics of various ground objects and modeling the global contextual relationships among them, which enhances the model’s robustness across diverse remote sensing applications.
+The LGWA module incorporates three local window attention mechanisms with varying sizes and configurations, alongside a single global attention mechanism.
+The distinct shapes of the local windows are specifically designed to extract different levels of local features.
+The horizontal window concentrates on capturing horizontally-oriented local details, the vertical window focuses on vertically-aligned local features, and the rectangular window is dedicated to extracting omnidirectional local information.
+This multi-faceted approach ensures the comprehensive extraction of multi-level features from ground objects.
+Subsequently, the global attention mechanism models the global contextual relationships among these objects based on their extracted local features, which enhances the overall performance of the model.
+Figure 4: Architecture of the LGWA-based Local Global Block, employing multiple local windows to extract local feature and single global window to extract global feature.
 For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@
 where $H_{i}=\text{Attn}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$, and $W^{O}\in\mathbb{R}^{hd_{v}\times d_{M}}$ projects the concatenated output back to the original dimension.
-Through this modular design, LGWA effectively captures both local structural details and global semantic context.
-The flexible design of window configurations allows the module to balance modeling power and computational efficiency, making it well-suited for multi-scale, high-resolution remote sensing applications.
 IV Experiments and Results
@@
 In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets.
-These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatio-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets.
-The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal semantic segmentation, binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
+These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatial-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets.
+The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
 TABLE I: Brief Introduction of The Experimental Datasets.
@@
 LoveDA Urban [40] LULC SS 1 1 3 7 1024×1024 0.3 5987
-Dynamic Earthnet [41] LULC SCD &SS &BCD 24 24 4 6 1024×1024 3 54750
+Dynamic Earthnet [41] LULC BCD &SCD 24 24 4 6 1024×1024 3 54750
 IV-A Datasets
@@
 To evaluate the effectiveness of the proposed STSUN, we conducted comparative experiments with various benchmark methods on building and LULC scene datasets.
-Since STSUN is adaptable to multiple datasets, it was trained and validated across all datasets in various scenarios, whereas the benchmark methods were trained and validated on individual datasets.
+Since STSUN is adaptable to multiple datasets, it was trained across all datasets in each scenario, whereas the benchmark methods were trained on individual datasets.
 On the four building scene datasets, the compared CNN-based models include FCN [42], SegNet [43], U-Net [44], PSPNet [45], HRNet [46], MA-FCN [47], Deeplabv3+ [48], ResUNet-a [49], MAPNet [50], D-LinkNet [51], SIINet [52], FC-EF [10], FC-Siam-Diff [10], FC-Siam-Conc [10], STANet [38], DTCDSCN [53], SNUNet [54], CDNet [55], DDCNN [56], DASNet [57], DSIFN [58], HANet [59], USSFCNet [60], and SEIFNet [61], the compared transformer-based models include Segformer [62], ChangeFormer [63] and A2Net [64], and the CNN-transformer hybrid models include BDTNet [65], TransUNet [66], CMTFNet [67], BIT [15], MTCNet [68], MSCANet [69], AMTNet-50 [70], Contrast-COUD [39] and TS-COUD [39].
-On the two LULC scene datasets, the compared CNN-based models include U-Net [44], Deepabv3+ [48], DANet [71], ResUNet-a [49], DASSN [72], HCANet [73], RAANet [74], A2-FPN [75] and CAC [76], the compared transformer-based models include LANet [77], SCAttNet [78], and TSViT [79], the CNN-transformer hybrid models include SAPNet [80], SSCBNet [81], UTAE [82], A2Net [83], SCanNet [84] and TSSCD [85].
+On the two LULC scene datasets, the compared CNN-based models include U-Net [44], Deepabv3+ [48], DANet [71], ResUNet-a [49], DASSN [72], HCANet [73], RAANet [74], MSAFNet [75], A2-FPN [76] and CAC [77], the compared transformer-based models include LANet [78], SCAttNet [79], CLCFormer [80] and TSViT [81], the CNN-transformer hybrid models include SAPNet [82], SSCBNet [83], UTAE [84], A2Net [85], SCanNet [86] and TSSCD [87].
 IV-C Implementation Details
@@
 In addition, following the settings of the DynamicEarthnet dataset [41], we use the semantic change segmentation (SCS) metric, classagnostic binary change score (BC) and semantic segmentation score among changed pixels (SC) metrics to evaluate the model’s performance on this dataset.
-Precision provides a measure of the incidence of false positives, while recall reflects the occurrence of false negatives.
-Due to the inherently inverse relationship between these two metrics, simultaneously securing high scores for both represents a significant challenge.
-The F1-score, calculated as the harmonic mean of precision and recall, offers a balanced assessment.
-The IoU metric evaluates spatial accuracy by quantifying the overlap between the predicted and ground truth changed pixels relative to their union.
-IV-D Ablation Study
-To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset.
-The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions.
-As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”).
-Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension.
-The DUM is then used to map the combined channel dimension into a unified feature.
-In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task.
-Concurrently, LGWA was substituted with standard global attention.
-This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
-TABLE II: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
-Strategy Attention F1(%) IoU(%) OA(%)
-Coupled Unification Global Attention 62.87 45.85 94.38
-Coupled Unification LGWA 63.22 46.22 94.72
-Decoupled Unification Global Attention 65.91 49.15 97.21
-Decoupled Unification LGWA 66.48 49.79 98.05
-Table II displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives.
-Furthermore, their combined use leads to even greater performance improvements.
-Specifically, the decoupled unification strategy enhances performance by preserving an independent temporal dimension.
-This allows STSUN to extract features from remote sensing images at various time steps during the encoder stage.
-Subsequently, in the feature fusion stage, feature-level fusion effectively reduces interference from irrelevant information [13].
-Moreover, when performing specific tasks, this strategy enables more effective integration with task embeddings and allows for the adjustment of temporal length to suit task requirements.
-This capability facilitates the joint modeling of multiple dense prediction tasks, leveraging data from various tasks to collectively improve their individual performance.
-In a similar vein, LGWA offers advantages over standard global self-attention.
-By employing a combination of variously shaped local windows alongside global self-attention mechanisms, LGWA is capable of concurrently extracting a rich tapestry of local features and comprehensive global context.
-This simultaneous extraction of multi-level features is crucial for enabling the model to effectively perform dense prediction tasks across various scales.
-IV-E Overall Comparison
-IV-E1 Building Scenario
-The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection.
-STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
+IV-D Overall Comparison on the Building Scenario
+The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection.
+STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table II, III, IV and V.
 Figure 5: Sample inference results on for building scene datasets.
@@
 On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed STSUN achieves state-of-the-art performance.
-As detailed in Table III, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%.
+As detailed in Table II, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%.
 This superior performance can be attributed to STSUN, particularly its spatial unification component, which effectively processes varying image size, and the Local-Global Window Attention mechanism.
 The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
-TABLE III: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
+TABLE II: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
 Methods P (%) R (%) F1 (%) IoU (%)
@@
 STSUN 95.71 94.87 95.29 91.00
-For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table IV.
+For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table III.
 The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points.
@@
 The STSUN component allows the network to effectively learn temporal evolutionary patterns and relationships between buildings and other objects, leading to more precise change maps.
-TABLE IV: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
+TABLE III: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
 Methods P (%) R (%) F1 (%) IoU (%)
@@
 In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, STSUN surpasses compared methods.
-Table V shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%.
+Table IV shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%.
 The strength of STSUN on this dataset lies in its robust temporal modeling capabilities, which allows for consistent feature representation across different time points.
 Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
-TABLE V: Accuracy Comparison on the LEVIR-CD Dataset. The Best Values Are Highlighted in Bold.
+TABLE IV: Accuracy Comparison on the LEVIR-CD Dataset. The Best Values Are Highlighted in Bold.
 Methods P (%) R (%) F1 (%) IoU (%)
@@
 STSUN 93.17 90.07 91.59 84.49
-Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%.
+Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table V, with an F1-score of 66.48% and an IoU of 49.79%.
 The TSCD dataset requires robust modeling of temporal dependencies across multiple observations.
 STSUN is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of buildings, which enables STSUN to accurately identify changes in complex, evolving landscapes.
-TABLE VI: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
+TABLE V: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
 Methods F1 (%) IoU (%) OA (%)
@@
 Visually, STSUN consistently produces accurate and complete segmentation and change maps.
-The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations.
+The results exhibit few false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations.
 For instance, in building segmentation on the WHU dataset (Figure 5(a)), STSUN generates sharp edges and complete building shapes.
@@
 This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of STSUN and LGWA, to learn highly discriminative features and effectively model contextual relationships across the spatial, temporal and spectral dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
-IV-E2 LULC Scenario
+IV-E Overall Comparison on the LULC Scenario
 The efficacy of the proposed Spatial-Temporal-Spectral Unified Network is rigorously evaluated against SOTA methods on two challenging LULC datasets: LoveDA for single-temporal land cover classification and DynamicEarthNet for multi-temporal land cover classification.
-On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII.
+On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VI.
 This leading performance can be attributed to STSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information.
 This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery.
-TABLE VII: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
+TABLE VI: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
 Methods F1-score per category (%) AF(%) OA(%) mIoU(%)
@@
 STSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
-For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VIII.
+For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VII.
 The success of STSUN on this dataset underscores the effectiveness of the STSUN in explicitly modeling the temporal dimension.
 By leveraging temporal metadata to ensure consistency across sequences of varying lengths, STSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
-TABLE VIII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.
+TABLE VII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.
 […]
@@
 TSSCD [87] 12.0 19.4 4.7 33.9
-STSUN 29.9 38.9 21.8 55.7
+STSUN 30.3 38.9 21.8 55.7
 Figure 6 presents the LULC semantic segmentation results of our proposed STSUN on the single-temporal LoveDA dataset (Figure 6 a) and the multi-temporal Dynamic EarthNet dataset (Figure 6 b).
@@
 (b) DynamicEarthNet dataset sample.
+IV-F Ablation Study
+To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset.
+The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions.
+As the temporal dimension is strongly correlated with dense prediction task types, this strategy enables joint modeling of multiple tasks on independent dimensions, learning the complementarity between each task.
+As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”).
+Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension.
+The DUM is then used to map the combined channel dimension into a unified feature.
+In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task.
+Concurrently, LGWA was substituted with standard global attention.
+This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
+TABLE VIII: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
+Strategy Attention F1(%) IoU(%) OA(%)
+Coupled Unification Global Attention 62.87 45.85 94.38
+Coupled Unification LGWA 63.22 46.22 94.72
+Decoupled Unification Global Attention 65.91 49.15 97.21
+Decoupled Unification LGWA 66.48 49.79 98.05
+Table VIII displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives.
+Furthermore, their combined use leads to even greater performance improvements.
+Specifically, the decoupled unification strategy enhances performance by preserving an independent temporal dimension.
+This allows STSUN to extract features from remote sensing images at various time steps during the encoder stage.
+Subsequently, in the feature fusion stage, feature-level fusion effectively reduces interference from irrelevant information [13].
+Moreover, when performing specific tasks, this strategy enables more effective integration with task embeddings and allows for the adjustment of temporal length to suit task requirements.
+This capability facilitates the joint modeling of multiple dense prediction tasks, leveraging data from various tasks to collectively improve their individual performance.
+In a similar vein, LGWA offers advantages over standard global self-attention.
+By employing a combination of variously shaped local windows alongside global self-attention mechanisms, LGWA is capable of concurrently extracting a rich tapestry of local features and comprehensive global context.
+This simultaneous extraction of multi-level features is crucial for enabling the model to effectively perform dense prediction tasks across various scales.
+IV-G Unification of Dense Prediction Tasks
+Semantic segmentation, binary change detection, and semantic change detection are prevalent dense prediction tasks in remote sensing, exhibiting significant similarities and complementarities.
+To validate the capability of STSUN to unify these diverse dense prediction tasks, we conducted a comparison experiment on four datasets in the building scenario.
+Specifically, we compare the performance of STSUN models trained individually on each of the four datasets against a single STSUN model trained jointly on all datasets.
+For a given dataset, the former is denoted as $STSUN_{\{dataset\}\_single}$ and the latter as $STSUN_{\{dataset\}\_unified}$.
+The results, presented in Table IX, demonstrate that $STSUN_{\{dataset\}\_unified}$ consistently outperforms its single-task counterparts across all building datasets.
+This performance gain is attributed to two factors.
+First, joint training on data from multiple dense prediction tasks exposes the model to a richer and more diverse set of samples, enhancing its ability to learn comprehensive spatial-temporal-spectral features from remote sensing imagery.
+Second, the model capitalizes on the complementary nature of these tasks.
+This synergy, analogous to a multi-task learning paradigm, enables the learning of more powerful and robust feature representations, thereby elevating the model’s performance across all individual tasks.
+TABLE IX: Comparison over Task Unification on Four Building Scenario Datasets. The Best Values Are Highlighted in Bold.
+Methods P (%) R (%) F1 (%) IoU (%)
+$STSUN_{whu\_single}$ 95.24 94.30 94.77 90.06
+$STSUN_{whu\_unified}$ 95.71 94.87 95.29 91.00
+$STSUN_{whucd\_single}$ 92.22 90.89 91.55 84.42
+$STSUN_{whucd\_unified}$ 93.07 91.20 92.13 85.40
+$STSUN_{levircd\_single}$ 92.94 89.63 91.25 83.92
+$STSUN_{levircd\_unified}$ 93.17 90.07 91.59 84.49
+$STSUN_{tscd\_single}$ 66.91 65.08 65.98 49.23
+$STSUN_{tscd\_unified}$ 67.32 65.67 66.48 49.79
+IV-H Dense Prediction with Flexible Semantic Category Set
+Remote sensing scenarios often encompass distinct sets of ground objects, necessitating different sets of semantic categories.
+For instance, the two LULC datasets employed in this study exhibit this variance: the LoveDA dataset uses the semantic set {building, road, water, barren, forest, agriculture, background}, whereas the DynamicEarthNet dataset uses {impervious surface, agriculture, forest, wetlands, soil, water}.
+To demonstrate STSUN’s ability to handle flexible semantic category sets for dense prediction, we designed an experiment using these two datasets.
+For this experiment, the DynamicEarthNet dataset was used exclusively for semantic segmentation, utilizing only the image pairs with semantic annotations.
+We compare four model configurations:
+  1. 1.
+    $STSUN_{loveda}$, trained exclusively on the LoveDA dataset to predict its 7-category set.
+  2. 2.
+    $STSUN_{dynamic.}$, trained exclusively on the DynamicEarthNet dataset to predict its 6-category set.
+  3. 3.
+    $STSUN_{fixed}$, trained jointly on both datasets to predict a fixed, 10-category set corresponding to the union of their individual category sets.
+  4. 4.
+    $STSUN_{flexible}$, trained jointly on both datasets but dynamically predicting from the appropriate category subset for each respective dataset.
+As shown in Table X, $STSUN_{flexible}$ outperforms $STSUN_{fixed}$ on both datasets, while achieving performance comparable to the specialist models, $STSUN_{loveda}$ and $STSUN_{dynamic.}$.
+The degraded performance of $STSUN_{fixed}$ is expected.
+This model is constrained to predict over the union of all categories, even for an image from a scene that does not contain certain categories.
+Forcing a single, fixed-size output space across scenes with disparate semantic sets introduces ambiguity and negatively impacts performance on each respective scene.
+In contrast, $STSUN_{flexible}$ dynamically adapts its predictive output to the relevant subset of semantic categories for each scene, which explains why its performance is on par with the individually trained models.
+Crucially, however, $STSUN_{flexible}$ offers superior model efficiency, as a single, unified model can be deployed across scenes with different semantic category sets without requiring any additional training or fine-tuning.
+TABLE X: Comparison over Category Unification on Two LULC Scenario Datasets. The Best Values Are Highlighted in Bold.
+$Model_{loveda}$ AF(%) OA(%) mIoU(%)
+$STSUN_{loveda}$ 74.66 71.70 65.59
+$STSUN_{fixed}$ 74.03 70.49 64.67
+$STSUN_{flexible}$ 74.81 71.82 65.73
+$Model_{dynamic.}$ SCS BC SC mIoU(%)
+$STSUN_{dynamic.}$ 30 38.4 21.6 55.3
+$STSUN_{fixed}$ 29.4 38.7 20.1 54.2
+$STSUN_{flexible}$ 30.3 38.9 21.8 55.7
 V DISCUSSION
-V-A Unified Multitask Learning
+V-A Unified Input and Output Modeling
+A core strength of the proposed STSUN framework lies in its capacity to accommodate arbitrary input and output configurations across the spatial, temporal, and spectral dimensions, overcoming the rigidity inherent in conventional architectures.
+This adaptability is particularly advantageous for processing the heterogeneous data prevalent in remote sensing applications.
+For input data, the spectral dimension $C_{1}$ is strongly correlated with sensor modality.
+The flexibility of STSUN in spectral dimension allows it to seamlessly ingest and model data from diverse sources, including RGB, Synthetic Aperture Radar, multispectral, and hyperspectral sensors, within a singular and unified model.
+At the same time, the temporal dimension $T_{1}$ is directly related to the time span and acquisition frequency of an image series.
+The ability to handle variable temporal lengths enables STSUN to process datasets with disparate temporal characteristics without modification.
+Moreover, the spatial dimensions $H_{1},W_{1}$ are tied to the geographic coverage and spatial resolution of the imagery.
+STSUN’s architectural design is agnostic to input image size, thus capably handling data from different regions coverage and spatial resolutions.
+This principle of unification extends symmetrically to the model’s output, enabling versatile and efficient inference.
+The configuration of the output temporal dimension $T_{2}$ is intrinsically linked to the dense prediction task being executed.
+This allows STSUN to unify semantic segmentation, binary change detection, and semantic change detection, facilitating joint modeling that can exploit complementary regularities between these tasks to enhance overall performance.
+Furthermore, the output channel dimension $C_{2}$ corresponds directly to the set of predicted semantic categories.
+By treating the category set as a flexible parameter, STSUN can adapt to diverse prediction class sets across various remote sensing scenarios.
+This obviates the need for extensive retraining or the maintenance of multiple specialized models, representing a significant step towards more scalable and universally applicable frameworks for remote sensing data analysis.
+V-B Unified Multitask Learning
 STSUN introduces a unified framework for dense prediction, capable of concurrently addressing semantic segmentation, binary change detection, and semantic change detection within a single model.
@@
 This results in a framework that is not only more data-efficient but also exhibits enhanced adaptability and robustness across a diverse range of dense prediction challenges.
-V-B Limitation and Expectation
-The increasing heterogeneity of remote sensing data across spatial, temporal, and spectral dimensions presents a fundamental challenge to conventional deep learning models, which are typically engineered for fixed data configurations and singular tasks.
-This rigidity impedes the integrated analysis of multi-source data, such as fusing Landsat and Sentinel imagery for longitudinal land cover analysis or adapting a single model for both routine monitoring and rapid disaster assessment.
-Our work addresses this critical gap by introducing a unified framework, STSUN, which demonstrates that a single model can adeptly navigate these STS variations.
-The scientific significance of this approach lies in its ability to holistically leverage the vast and disparate archives of remote sensing data, fostering the development of more robust, flexible, and data-efficient solutions for dense prediction.
-Our experimental results robustly demonstrate that a single STSUN can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks.
-Across multiple datasets, STSUN consistently achieved performance comparable to or exceeding that of state-of-the-art specialized models, highlighting its significant potential for streamlining complex remote sensing data analysis workflows and reducing the need for task-specific model development.
-The superior performance of STSUN, as evidenced by our extensive experiments, stems from its departure from traditional multi-task learning paradigms.
-Prevailing MTL methods in remote sensing often require co-registered, fully labeled datasets for all tasks and rely on multiple task-specific decoders.
-In contrast, STSUN’s architecture, centered on a shared decoder modulated by task and category embeddings, facilitates joint modeling on a diverse corpus of non-paired data aggregated from semantic segmentation, binary change detection, and semantic change detection.
-Training on this wider spectrum of data forces the model to learn a more generalized and powerful feature representation by extracting shared knowledge across tasks and data sources, leading to enhanced performance without the need for task-specific architectural components.
-Despite its promising results, the current implementation of STSUN has two principal limitations that suggest clear avenues for future research.
-First, the framework’s flexibility is contingent upon explicit metadata, such as predefined task and category embeddings, which must be provided at inference time.
-A compelling future direction involves exploring methods for the model to implicitly infer the task and desired output configuration directly from the input data context, thereby increasing its autonomy.
-Second, the potential of our unified modeling approach has been demonstrated only within a supervised learning context.
-A significant opportunity exists to extend this framework to the self-supervised domain to construct vision foundation models and vision language models.
-By pre-training the vision model on massive, unlabeled remote sensing archives with diverse STS characteristics, one could develop a powerful remote sensing vision foundation model.
-Such a model, pre-trained to understand the fundamental structure of STS data, could then be fine-tuned with minimal data for a wide array of downstream applications or potentially used in zero-shot settings, drastically reducing the dependence on labeled data and advancing the capabilities of large-scale Earth observation.
-At the same time, applying this method to remote sensing vision-language models could enable the models to be trained with large amounts of remote sensing image-text pair data without needing to consider the dimensionality settings of the remote sensing images.
-This would not only enhance the performance of remote sensing vision-language models but also make them adaptable to various STS settings for remote sensing image and text pair data.
+V-C Limitations and Expectations
+While the proposed Spatial-Temporal-Spectral Unified Network demonstrates considerable promise in harmonizing the analysis of heterogeneous remote sensing data, its current instantiation presents certain limitations which, in turn, illuminate clear and compelling trajectories for future research.
+First, the operational flexibility of our framework is presently contingent upon the provision of explicit metadata at inference time, specifically in the form of predefined task and category embeddings.
+This requirement, while effective, curtails the model’s autonomy and presupposes a level of a priori knowledge about the analytical objective.
+A significant advancement would be to imbue the model with the capacity to implicitly infer the task and desired output configuration directly from the contextual cues within the input data stream.
+Future work could explore methodologies grounded in meta-learning or employ sophisticated attention mechanisms that learn to dynamically weigh different aspects of the data, thereby deducing the analytical intent without explicit instruction.
+Second, while our method achieves a critical unification at the data format and architectural input level across the spatial, temporal, and spectral dimensions, this does not inherently guarantee robust semantic generalization across the vast heterogeneity of remote sensing scenes, sensor modalities, and non-training semantic category sets.
+A promising path to surmount this limitation lies in elevating the STSUN framework into a large-scale, foundational model for Earth observation by employing advanced self-supervised or multi-modal pre-training strategies.
+The objective would be to produce a model that captures the fundamental structures of spatial-temporal-spectral data, enabling highly transferable representations that dramatically improve performance on a wide array of tasks with minimal, or even zero-shot, fine-tuning.
+Finally, the potential of our unified modeling paradigm has thus far been demonstrated exclusively within a supervised learning context.
+A significant opportunity exists to extend this framework into the self-supervised and vision-language domains.
+By pre-training the unified vision model on enormous, unlabeled remote sensing archives, one could construct a powerful Remote Sensing Vision Foundation Model.
+Such a model, pre-trained to comprehend the elemental structure of diverse STS data, could then be adapted with remarkable data efficiency for specialized downstream tasks.
+Concurrently, applying this unified input approach to Remote Sensing Vision-Language Models holds transformative potential.
+It would permit the training of these models on vast corpora of image-text pairs without the need for cumbersome, modality-specific engineering to handle varying spectral or temporal dimensions in the imagery.
+This would not only streamline the development and enhance the performance of RS-VLMs but also make them natively adaptable to the full diversity of remote sensing data, fostering a new generation of models capable of nuanced, cross-modal understanding of our planet.
 VI CONCLUSION
@@
 STSUN uniquely integrates semantic segmentation, binary change detection, and semantic change detection within a single model via trainable task embeddings, obviating the need for specialized, task-specific architectures.
-Moreover, STSUN utilizes trainable semantic category embeddings to perform dense prediction for any prediction class setting without requiring model retraining.
+Moreover, STSUN utilizes trainable semantic category embeddings to perform dense prediction for multiple prediction class setting without requiring model retraining.
 Comprehensive experimental validation on multiple datasets confirmed that a single STSUN model consistently adapts to varied data inputs and outputs, unifing multiple dense prediction tasks and prediction category settings, achieving or exceeding state-of-the-art performance.